### v2

# ConCap: Practical Network Traffic Generation for (ML- and) Flow-based Intrusion Detection Systems

Abstract
Network Intrusion Detection Systems (NIDS) have been studied in research for almost four decades. Yet, despite thousands of papers claiming scientific advances, a non-negligible number of recent works suggest that the findings of prior literature may be questionable. At the root of such a disagreement is the well-known challenge of obtaining data representative of a real-world network—and, hence, usable for security assessments.
We tackle such a challenge in this paper. We propose ConCap, a practical tool meant to facilitate experimental research on NIDS. Through ConCap, a researcher can set up an isolated and lightweight network environment and configure it to produce network-related data, such as packets or NetFlows, that are automatically labeled—hence ready for fine-grained experiments. ConCap is rooted on open-source software and is designed to foster experimental reproducibility across the scientific community by sharing just one configuration file. Through comprehensive experiments on 10 different network activities, further expanded via in-depth analyses of 21 variants of two specific activities and of 100 repetitions of four other ones, we empirically verify that ConCap produces network data resembling that of a real-world network. We also carry out experiments on well-known benchmark datasets as well as on a real “smart-home” network, showing that, from a cyber-detection viewpoint, ConCap’s automatically-labeled NetFlows are functionally equivalent to those collected in other environments. Finally, we show that ConCap enables to safely reproduce sophisticated attack chains (e.g., to test/enhance existing NIDS). Altogether, ConCap is a solution to the “data problem” that is plaguing NIDS research.
I Introduction
There is an anomaly in research on Network Intrusion Detection Systems (NIDS). On the one hand, a deluge of papers continues to expand the boundaries of our knowledge, as shown by khraisat2019survey ; apruzzese2023sok ; ahmad2021network . On the other hand, a growing number of recent efforts highlight some “pitfalls” that undermine the foundations of prior research (e.g., arp2022and ; engelen2021troubleshooting ; liu2022error ; catillo2023machine ; flood2024bad ).
To portray this contrast, we can look at the panorama of open-source datasets used in NIDS-related research, some of which count thousands of citations according to Google Scholar flood2024bad . As a concrete example, consider the paper presenting the CICIDS17 dataset sharafaldin2018toward : published in 2018, this paper had 1600 citations in Q3 2022, which increased to 4200 in Q4 2024—indicating a substantial growth in NIDS research. Yet, in 2021, Engelen et al. engelen2021troubleshooting pinpointed glaring issues in CICIDS17—particularly in terms of ground-truth labeling of attack samples. The flaws of CICIDS17 (and also of its successor, CICIDS18) have been “fixed” in 2022 liu2022error . However, the EuroS&P’24 Best Paper Award flood2024bad revealed that most datasets used by prior research have “bad design smells.”
Simply put, the stark reality is that (i) data is required to support a paper’s claims, but (ii) high-quality data is hard to come by in the NIDS context—especially from the viewpoint of an academic researcher apruzzese2023sok ; flood2024bad ; cordero2021generating . To aggravate this problem, modern NIDS increasingly rely on data-driven techniques, such as machine learning (ML). Therefore, labeled data is necessary to properly evaluate the pros and cons of state-of-the-art NIDS apruzzese2023sok . Finally, despite the benefits that open-source datasets (under the assumption that they are correctly labeled) can provide to a researcher bonninghausen2024introducing , exclusive reliance on such “static benchmarks” prevents one from generating new data. This impairs the assessment of existing NIDS against recent, and more sophisticated threats—such as multi-step attacks navarro2018systematic ; mitreattack , or attacks exploiting recently-discovered vulnerabilities cve . We aim to rectify such a “data problem.”
Our major technical contribution is ConCap, an open-source system to generate (and label) network-traffic data mimicking that of a real-world network. ConCap is particularly suited to generate network data pertaining to malicious activities. Such data can then be used alongside “benign” data taken from the network environment that the NIDS is meant to protect—which is a well-founded assumption arnaldo2019holy ; apruzzese2022cross ; cordero2021generating . Practically, such environment can be: (a) that of a benchmark dataset bonninghausen2024introducing , or of (b) an ad-hoc network testbed for experimental research gomez2023survey , or even that of (c) a real-world network apruzzese2022role . We design ConCap so that it is (i) flexible enough to enable reproduction of any (allegedly malicious) network activity, whose generated datapoints are (ii) automatically labeled at the granular level, while also enabling (iii) control of the network conditions (e.g., to simulate resource starvation). As such, ConCap represents a solid foundation to address the “data problem” that affects NIDS research. We will demonstrate this.
To develop ConCap, we assembled open-source technologies, such as container-related frameworks, networking utilities, and network flows (NetFlow apruzzese2023sok ; verkerken2025rustiflow ) extractors. Our design facilitates experimental reproducibility: to (automatically) generate labeled NetFlows pertaining to one (or more) attacks, ConCap only requires the researcher to define a “scenario,” i.e., a file describing the activities of the “attacker” and “target” host(s), as well as the characteristics of the overarching network channel. In this way, other researchers can replicate the same experiments just by running the same “scenario” in their own version of ConCap, thereby removing the need to share data—which can be quite big in size (e.g., the CICIDS18 sharafaldin2018toward dataset is larger than 400GB). Such a property intrinsically fosters scientific reproducibility—which is lacking even in top-tier security venues olszewski2023get ; apruzzese2023sok .
We empirically validate the realism of ConCap’s generated data. Through extensive experiments wherein we compare the network traffic produced by a physical bare-metal environment with that generated by ConCap, we find that there are negligible differences (at both the packet- and NetFlow-level) between these two setups. We deeply investigate such differences, and confirm that they are due to the well-known “unpredictable” behavior of modern networks sommer2010outside ; apruzzese2023sok . Inspired by such a finding, we go one step further and examine if the data generated by ConCap is deterministic—an aspect that received limited attention from related research (e.g., clausen2019traffic ). Our analysis, entailing 100 executions of the same set of four distinct network activities, reveals that ConCap’s output cannot be claimed to be fully deterministic—as is the case for real-world networks, too. However, through other experiments on well-known benchmarks and on a real “smart-home” network, we show that, ML-wise, the data generated by ConCap is functionally equivalent to that of other network setups.
Altogether, our evaluations demonstrate that ConCap represents a convenient tool to address the “data problem” that affects NIDS-related research. For instance, we also show how to use ConCap to carry out security assessments of NetFlow- and ML-based NIDS against recent CVE; as well as how to replicate a sophisticated attack chain, resembling nine MITRE ATT&CK tactics, and generate the corresponding (labeled) traffic. Finally, we also show that ConCap is lightweight: it takes less than 3 seconds to power-up on a laptop, and its memory footprint takes less than 40MB.
Contributions. After positioning our paper within extant literature and motivating our work (in Section §II), we:
  • Propose ConCap, a system for generating ad-hoc network traffic (§III), particularly suited to create and label malicious datapoints, facilitating network security research; we extensively describe and justify our implementation choices—rooted in open-source software and reproducibility (§IV).
  • Empirically demonstrate (§V) that ConCap generates network-traffic data resembling that of bare-metal setups, and prove that experiments run via ConCap are not fully deterministic—a behavior expected by real-world networks.
  • Show that (§VI), from an ML viewpoint, the (labeled) NetFlows produced via ConCap are functionally equivalent to those originating from other setups—such as from benchmark datasets, or from a real “smart home” network. We also show that ConCap can generate data of sophisticated multi-step attacks and recent threats to test existing NIDS.
We also discuss our findings (§VII) and compare ConCap with related works (§VIII). We release all our resources repository .
II Related Work and Motivation
We summarize the domain of NIDS (§II-A) and we outline the challenges of acquiring data for assessing NIDS (§II-B), representing the root of the problem tackled by our paper. Then, we motivate our technical (§II-C) contribution, which specifically focuses on NIDS analysing network flows (NetFlows vormayr2020my ) due to the widespread usage of this datatype apruzzese2023sok ; flood2024bad ; dias2020go .
II-A Network Intrusion Detection Systems (NIDS)
Modern networks are constantly under attack enisa2023threat ; cloudflare2024threat , and NIDS represent the first line of defense against the ever-evolving cyberthreats sans2024report . The main goal of an NIDS [^1] Borrowing from apruzzese2023sok , we use “NIDS” in a broad sense, including also SIEM comodo2023idssiem or EDR paloalto2023edrsiem (which can be seen as extensions of NIDS). is to detect any given threat (e.g., a botnet-infected host, a remote DDoS attack, or an attacker who acquired access to an internal host) as early as possible so that proper mitigations can be enacted to mitigate the damage of an offensive campaign bertino2021ai . Fig. 1 shows a schematic of an NIDS deployment scenario.
Figure 1: Exemplary deployment scenario of an NIDS.
The advent of data-driven technologies, such as AI/ML (which support both signature- and anomaly-based methods apruzzese2023sok ), has been adopted by NIDS developers kshetri2021economics . Yet, even state-of-the-art NIDSs struggle with the sheer amount of attacks that target current organizations enisa2023threat . Intriguingly, even though practitioners do appreciate the analytical capabilities of ML mink2023everybody , modern security operation centers are overwhelmed with false alarms alahmadi202299 ; vermeer2023alert . Put simply, despite being an instrumental security tool, there is a constant need to improve NIDS—creating a fertile ground for research.
Since the seminal work by Denning denning1987intrusion , thousands of scientific papers have sought to propose (e.g., mirsky2018kitsune ; barradas2021flowlens ; piskozub2021malphase ; fu2021realtime ), enhance (e.g., sommer2003enhancing ; araujo2019improving ), or assess (e.g., apruzzese2022cross ; jacobs2022ai ; corona2013adversarial ) NIDS. Despite all such efforts, a recent SoK apruzzese2023sok revealed that there is skepticism among industry experts with regard to the results claimed in research. Such doubts are well-founded: various recent papers (e.g., liu2022error ; catillo2023machine ; flood2024bad ) identified critical issues that are becoming endemic in research. The root cause of most such critiques is the poor quality of the data used to test these systems—which is a problem that has been known to affect this research domain since at least 2010 sommer2010outside .
II-B Research Challenge: Obtaining Data for NIDS
Any security tool must be tested before its deployment. Such tests require data. For NIDS, the evaluation should be carried out on data that is representative of the environment in which the NIDS is meant to be deployed apruzzese2023sok . Unfortunately, carrying out the abovementioned operations is tough from the perspective of an (academic) researcher. Let us outline the options that enable one to collect network-related data for a security assessment of a NIDS, explaining their challenges.
  • Real-world capture from the deployment environment. This is ideally the best option since it guarantees that the data resembles the one that will be generated by the monitored network. However, such an option may not be available: the researcher may not have access to such data due to privacy reasons, and infecting/attacking physical devices may not be acceptable in some organisations (even if for research wangr2025rr ).
  • Synthetic capture from a custom environment. This is a sensible alternative: by creating an ad-hoc network (e.g., via virtual machines sharafaldin2018toward ), a researcher has plenty of freedom to collect and generate any sort of data. However, the benign data may not be representative of the deployment environment. Moreover, even in such a setup, it is currently challenging to precisely distinguish benign from malicious data points: as shown in liu2022error ; flood2024bad , there is a risk of mistakenly “label” benign samples as malicious. Such errors can skew the results of the final assessment krauss2024verify .
  • Reliance on benchmark datasets. The last option is to use publicly available data, e.g., generated by other researchers in their own environments (either from physical or virtualized setups). This is a convenient option: it is exempt from privacy concerns and requires minimal technical expertise (since no simulation occurs). However, the validity of the corresponding evaluation will depend on whether the benchmark is a meaningful representation of the network wherein the NIDS is meant to be deployed.
For real-world deployments, it is paramount to evaluate the NIDS in the (real) network to be monitored by the NIDS: as highlighted by Sommer and Paxson sommer2010outside , networks present immense variability. Hence, even if any given NIDS is shown to “work well” on data from a custom network, it is questionable whether the same NIDS works well also in other networks. This is known as the “generalisability” verkerken2022towards ; sarhan2022evaluating (or “transferability” catillo2022transferability ) problem of NIDS, which prevents the creation of plug-and-play NIDS (confirmed by practitioners apruzzese2022role ). However, a research paper needs not to aim for real-world deployment to provide a significant contribution to the state of the art apruzzese2023sok . Our work is rooted in this truth: we focus on improving future research on NIDS—which not necessarily requires assessments on “real” networks to be valuable.
II-C Practical Generation of Network Data
Prior research on NIDS suffers from a “data” problem. Our main goal is to provide a solution to this problem by enabling future research to carry out meaningful assessments of NIDS.
To elucidate the importance of our major contribution, we present a motivational example. Suppose a researcher has no access to a real-world network for NIDS assessments. How can such a researcher conduct meaningful experiments? From our prior descriptions (§II-B) we identify three possible options.
  • The researcher can create a simulated/virtual network, but doing so requires dealing with the labeling issues (for the malicious traffic), and is (likely) limited to small-scale evaluations due to the impossibility of recreating a large network environment clausen2019traffic .
  • The researcher can exclusively rely on benchmark datasets, but this would limit the evaluation to the data within the benchmark dataset (e.g., even by manipulating the benchmark’s datapoints, there is a risk of breaking domain constraints sheatsley2021robustness ), preventing exploration of new threats.
  • To overcome the abovementioned limitations, the researcher can do a mix of the above arnaldo2019holy ; apruzzese2022cross : they can use “benign” data collected from a public dataset (potentially validated by prior work liu2022error ), and then generate “malicious” data to use alongside the benign data to carry out a proper evaluation.
We argue that the third option is the most enticing one. As of 2025, there are various publicly available datasets captured in large networks (see bonninghausen2024introducing for a list) whose data can be treated as “benign” for experimental purposes. [^2] Such “benign” data can also come from the specific (real) network in which the NIDS is to be deployed—but, as we argued, this is not necessary (although it would increase the soundness of an evaluation). Hence, if one could generate “malicious” data so that (i) it is “correctly labeled,” and (ii) it resembles the data generated by the host of the “benign” network, then one would be able to test a NIDS against a wide array of cyber threats—including new ones.
Research Goal. We seek to develop an open-source tool that enables the automatic generation (and labeling) of network traffic data that resembles that of a real network.
Figure 2: Overview of ConCap. [left] ConCap configured with two NetFlow extractors and three scenarios, [mid] executing all scenarios simultaneously on the cluster. [right] A view of a running scenario’s attacker and target pods.
III Our Proposed Tool: ConCap
We present ConCap (short for “Container Capture”), our tool to generate realistic network traffic for research in NIDS. After stating the design objectives (§III-A) and core principles (§III-B) of ConCap, we summarize its workflow (§III-C), schematically depicted in Fig. 2. We will discuss the low-level implementation of ConCap in the next section (§IV).
III-A Objectives
The underlying purpose of ConCap is to simplify the operations involved in the generation and labeling of malicious network activities that resemble those in a physical and real-world network. ConCap seeks to fulfill three objectives:
  • Attack flexibility. The tool must allow one to specify and reproduce a variety of actions that lead to network communications—including those denoting “malicious” behavior. The tool shall support both “simple” use cases (e.g., a machine attacking a different machine) as well as more “sophisticated” ones (e.g., a complex offensive campaign envisioning multi-step attacks across multiple hosts).
  • Data collection and labeling. The tool must automatically: (i) capture network traffic related to the specified actions; (ii) generate statistical metadata summarizing the communications, i.e., NetFlows; (iii) assign a label to these NetFlows.
  • Parallelism and control. The tool must enable simultaneous execution of diverse experiments, and fine-grained control of the network conditions of each experiment—so to reproduce the behavior of any network, and enable “bulk” experiments.
To align these objectives with our overarching research goal (see §II-C), the entirety of ConCap’s source code (which must leverage open-source libraries/frameworks) is publicly released repository ; moreover, the authors will make an effort to ensure long-term maintenance of ConCap’s code repository.
III-B Core Principles
We outline the core principles of ConCap, showing that our three objectives are embedded in its underlying design.
ConCap is rooted on the concept of “scenario”, which serves as a customizable blueprint for an entire experiment. When defining a scenario, developers can specify: (i) the activities carried out by the involved hosts; (ii) the conditions of the overarching network environment; (iii) the fine-grained label to assign to the generated network traffic metadata.
ConCap executes each scenario in an isolated “containerized” environment that mimics the behavior of real, physical hosts, and allows for unlimited parallel execution of scenarios—subject to the resource constraints of the experimental testbed on which ConCap is executed.
A scenario in ConCap assumes a set of hosts: one designated as the “attacker”, and one or more designated as the “target”. These hosts are fully controllable by the developer through customizable container configurations and attack commands. ConCap captures the network packets (as a PCAP trace) exchanged between the attacker and the target(s), automatically extracts high-level NetFlow records, and then labels such NetFlows according to the scenario definition. Labeling accuracy is guaranteed by configuring the attacker host to only reproduce “malicious” actions, since the isolated environment ensures no background noise is present.
ConCap supports reproducing both simple attacks, as well as complex, multi-step attack chains. Such support is provided via the “‘scenario” configuration file. In simple-attack scenarios, the attacker host executes a sequence of actions against a single host; for instance, the scenario in Listing 1 (in the Appendix B) shows a simple port-scan done by one “attacker” host against one “target” host. In complex-attack scenarios, however, it is possible to designate multiple “target” hosts—which can be either used to reproduce an attack against multiple hosts (e.g., an horizontal port scan; or a multi-chain attack in which each step requires the completion of the previous steps). The network traffic of each target is captured individually and labeled according to the specifications provided in the scenario file, enabling fine-grained analysis.
III-C Architecture and Workflow
We present in Fig. 2 a schematic of a typical setup of ConCap, highlighting the most relevant logical units. We use Fig. 2 to explain the operational workflow and architecture of ConCap.
Input. On the left, the “scenarios” folder contains three YAML configuration files, each defining a given scenario to be executed by ConCap. An example scenario configuration is shown in Listing 1, highlighting the various options (i.e., attacker and target definition, network conditions, and labeling) that can be configured before executing any given experiment. For added flexibility, ConCap supports using different NetFlow generation tools. This can be specified through configuration files located in the “processingpods” folder, where the details of the desired tool are defined.
Execution. When executing a scenario, ConCap parses the configuration files and interacts with both the host machine and the cluster, consisting of one or more nodes. If multiple scenarios are specified, ConCap supports concurrent execution by distributing them across different nodes, enabling parallelism (see the central section of Fig. 2). ConCap follows the instructions in the scenario files to set up the attacker and target pods, apply custom network configurations, initiate traffic captures, trigger attack execution, and run processing pods for feature extraction and labeling. See the rightmost section of Fig. 2 for a detailed view of the attacker and a target pod during scenario execution.
Output. For each scenario, ConCap creates a dedicated folder containing all experiment outputs, including logs, the PCAP trace, and the labeled NetFlows (see left of Fig. 2). The reason why we focus (also) on NetFlows for ConCap is due to their widespread popularity for network-related experiments (both in research and in practice apruzzese2023sok ; oh2022deepcoffea ); for instance, most public benchmark datasets are also released in this format bonninghausen2024introducing . Such a design choice serves to facilitate future research.
IV Creation and Functionalities of ConCap
We first explain and justify the key design choices followed to create ConCap (§IV-A). Then, we describe the “scenario file,” which embeds the most original contributions of ConCap (§IV-B). Finally, we explain how ConCap enables also the simulation of sophisticated malicious activities, such as multi-step and multi-target attacks (§IV-C).
IV-A Design and Implementation Choices
We present and justify the elements that compose ConCap.
Isolated Containerized Environment. As explained (in §II-B), the major advantage of running network simulations is the ability to generate malicious network traffic without putting real-world networks at risk. Such simulations can leverage, e.g., virtual machines (as done, e.g., in sharafaldin2018toward ) or containers. Containers are a lightweight alternative to virtual machines potdar2020performance , which are becoming very popular in related research (e.g., clausen2019traffic ; catillo2024towards ; boettiger2015introduction ) also for the ease of reproducibility moreau2023containers and reduced resource requirements potdar2020performance . For this reason, we used containers to develop ConCap. There exist many open-source solutions that can be used to deploy containerized applications casalicchio2020state . For ConCap, we rely on Kubernetes kubernetes , due to its widespread adoption (>>110k GitHub stars kubernetesGithub ) and key advantages over alternatives, such as enabling deployment and management of workflows across multiple machines bentaleb2022containerization —which is required to achieve our design goals of simultaneous execution of scenarios. For comparison, DetGen clausen2019traffic relies on Docker docker , which by default does not support parallel scenario execution across multiple machines (we will compare ConCap with DetGen in §VIII-B). We stress, however, that Kubernetes alone does not provide the functionalities provided by ConCap (see §III-C): we simply use Kubernetes as the backbone.
Attacker and Target(s). In Kubernetes, a “pod” is the smallest deployable unit, functioning as a logical host that groups one or more containers with shared storage and networking resources. To ensure isolation and enable fine-grained control over each host involved in the “attack”, we create ConCap so that the attacker and target(s) hosts are deployed in separate pods. This allows to configure parameters such as bandwidth and latency for each host. The attacker pod runs one container that simulates the behavior of the attacker’s host. In contrast, each target pod runs two containers: one serving as the host targeted by the attacker, and another that captures the attacker-target network communications (via tcpdump, which can capture all traffic of the target host since containers within the same pod share networking resources).
NetFlow format. ConCap supports any type of NetFlow generation software, including novel ones (e.g., verkerken2025rustiflow ). Choosing a given NetFlow tool is done by editing a dedicated configuration file (shown in Listing 2 in the Appendix B). In our proposed implementation of ConCap, we have integrated CICFlowMeter and Argus, which are popular in research and open source vormayr2020my ; apruzzese2023sok . Recent surveys apruzzese2023sok ; verkerken2025rustiflow revealed that most papers on ML-NIDS recently published in various (including top-tier) security venues rely on NetFlow for their analyses, motivating our focus on NetFlow.
Reproducibility. An implicit objective of ConCap, from a scientific viewpoint, is to facilitate the reproducibility of network traffic generation experiments. Many existing public network traffic datasets lack detailed metadata/instructions about how the traffic was produced, often providing only high-level descriptions of the experimental setup flood2024bad . This lack of transparency makes it difficult to perform an in-depth analysis of such datasets and prevents researchers from linking individual network traffic to their specific causes or originators, which is critical for tasks such as automated detailed labeling. Reproducibility in this context means not only re-running the same scenario and obtaining consistent traffic captures and labels, but also sharing complete experimental configurations in a way that others can reliably replicate the results on different systems. To address this, ConCap encapsulates all aspects of the traffic generation experiment in a reusable scenario file. These files define the behavior of the attacker and target(s), their environment, and the network conditions, serving as a blueprint of the experiment. By versioning and sharing these files alongside the resulting datasets, ConCap enables other researchers to reproduce the traffic generation process with minimal manual setup and to understand the context behind the captured network traffic. Moreover, the framework enforces consistency by automating the full execution pipeline–from deployment and traffic capture to flow extraction and labeling–ensuring that experiments can be rerun with similar results.
IV-B Customisability (Scenario definition)
Among ConCap’s greatest advantages is its customisability—which we claim as our major original contribution. Indeed, we are not aware of any tool which provides the same degree of flexibility (we discuss related work in §VIII). Such advantages are enabled by our custom implementation of the scenario file, for which we have provided a snippet in Listing 1. In what follows, we provide more details on the functionalities provided by the scenario file. Recall that the scenario file follows a YAML notation which requires to specify four components: the attacker, the target, the network, and the label.
Host configuration. The components describing the “attacker” and “target” hosts require a name and the containerImage to deploy the containers that simulate their behavior. Additionally, we designed ConCap so that it is possible to specify the conditions of the computing runtime (e.g., CPU and memory): this is crucial for attacks that disrupt the availability of the target host (e.g., DoS). The “attacker” component also has the atkCommand (used to start the attack), and an optional atkTime parameter that controls the attack duration (in seconds): intuitively if $atkTime>0$, then the attack will stop after the provided number of seconds; otherwise, the attack will continue until it naturally terminates. We also allow to configure the startup probe of the “target” component, which determines when the target is ready, as well as the filter used by tcpdump for the packet capture (PCAP). Additionally, the attacker and/or target(s) can be run in privileged mode. This is useful for tools or services that need elevated permissions (e.g., raw socket access) or to simulate sophisticated attacks.
Networking. The network component defines the networking environment, enabling to specify, e.g., the bandwidth, latency, and packet loss of the communication channel. Such a functionality is useful to, e.g., simulate various conditions—not only for covering a wide array of “benign” network environments, but also to reproduce circumstances which can conceal traces of malicious activities (e.g., catillo2024towards ). We are not aware of any open-source traffic-generation tool that enables (by default) specific configurations of the network conditions at the host level. This component, which we implemented via the open-source Traffic Control tc , also allows to determine if the generated traffic follows a specific distribution.
Labeling. The label component describes the labeling logic applied to the NetFlows (generated after processing the PCAP file). This can be configured globally, and individually per host. Given that ConCap enables precise control of the entire attack workflow (i.e., attacker and target(s) host and network conditions), the assigned labels ensure the resulting NetFlows are associated with the correct ground truth (since they all share the same “malicious” generative process). Moreover, since certain cyberattacks can be part of more sophisticated offensive campaigns (e.g., a port-scan can be part of a lateral-movement operation apruzzese2017detection ), and since such use-cases are supported in ConCap, it is possible to assign more granular labels. For instance, in Listing 1 the “category” is “port-scan” whereas the subcategory is “nmap”: to better identify the corresponding malicious traffic in a lateral-movement context, it would have been possible to specify “lateral-movement” as “category”, and “port-scan” as “subcategory”.
IV-C Sophisticated attacks (multi-step and multi-host)
In its most simple form, ConCap can be used to generate network traffic entailing one attacker and one target host. However, we designed ConCap so to also enable simulation of more complex network activities, such as multi-host attacks (e.g., an attacker host seeking to scan multiple target hosts) or multi-step attacks (e.g., an attacker that seeks to brute force an SSH server after finding out that there is an active SSH service running on such a host) in a reproducible and structured way. Here, we describe how we enabled support of multi-step/host attacks in ConCap (which we also claim as an original technical contribution).
Challenges. Simulating such advanced scenarios is non-trivial. Multi-step attacks often involve transferring state between phases (e.g., discovered credentials apruzzese2017detection ) combined with dynamic network reconfiguration (e.g., proxy routing or port forwarding angelini2019mad ). Additionally, these attacks may require accessing internal services only reachable after a successful compromise of a given gateway, demanding precise orchestration of timing, sequencing, and conditional logic.
Solutions. To address these challenges, ConCap introduces several mechanisms. Scenario definitions support dynamic environment variable substitution—instead of being constrained by hard-coded variables. Moreover, ConCap’s built-in startup-readiness probes ensure that attacks are launched only when the targets are fully operational. Scenario-wide and per-target labels are automatically merged to annotate captured flows based on their context within the attack chain. Fig. 3 illustrates the fully-automated process enabled by ConCap. On the left, the developer defines an arbitrary and flexible attack chain using the “atkCommand” in the scenario file that specifies the attacker’s behavior. In the middle, ConCap automatically deploys the pods, applies the network configuration, and manages pod lifecycles and readiness. On the right, it performs per-target traffic capture, flow extraction, and labeling. This workflow is repeated for each custom-defined step of the attack chain. In a sense, ConCap acts as a supervisor guaranteeing that each step is executed at the right time, thereby enabling smooth simulation of multi-step attacks—potentially requiring information acquired after intermediate steps. An end-to-end demonstration of using ConCap to reproduce a multi-step/multi-target attack chain, resembling acknowledged MITRE ATT&CK tactics, is provided in the Appendix D-B.
Figure 3: Multi-target scenario with ConCap. [left] ConCap supports advanced multi-step attack chains, [mid] performed over multiple targets. [right] The traffic is captured and processed per target, enabling per-target labeling.
Takeaway. ConCap is the first open-source tool for generating realistic, labeled network traffic for NIDS research. It offers attack flexibility, automated NetFlow generation and labeling, and parallel scenario execution with fine-grained control. A short demo is available in our repository repository .
V Real-world Validation of ConCap
The design choices we followed to develop ConCap (e.g., the usage of containers, or the usage of Traffic Control) indicate that ConCap is likely to generate traffic that mimics that of a physical real-world network. In what follows, we empirically confirm this hypothesis—a validation that has not been done in most prior works proposing traffic-generation tools (e.g., zhoutrafficformer ). Specifically, we ask ourself our first research question (RQ1):
Research Question #1: Does ConCap generate network traffic that resembles that of a physical network?
Motivation. For instance, while implementing ConCap, we may have made some mistakes; or it is possible that the usage of containers may lead to different traffic (w.r.t. that of a physical network). It could also be that there are some intrinsic differences: in which case, it is instructive to understand their root cause. Indeed, we are not aware of prior works on containers empirically testing such an hypothesis. Hence, our assessment serves to gauge whether our implementation of ConCap is correct, and that ConCap meets our overarching research goal (see §II-C); as well as studying potentially unknown properties of containers from a network perspective.
Approach. To answer RQ1, we carry out an intuitive set of experiments: we compare the network data generated by ConCap to that of a physical “bare-metal” setup—across a wide range of network activities. Such comparisons entail both quantitative and qualitative assessments, both at the network-packet level, as well as at the network-flow level. In particular, to comprehensively address RQ1, we perform three different experiments. First, as a starting point, we execute ten network activities, and see if there are any substantial differences from a quantitative viewpoint (§V-A). Second, we conduct an in-depth analysis focused on examining over twenty different variants of two specific activities types (§V-B). Third, as an ancillary experiment, we study the extent to which the traffic generated by ConCap can be considered to be deterministic (§V-C).
Common Setup and Workflow. The experiments discussed in this section entail two distinct environments.
  • Physical network. Two “bare-metal” physical hosts, each having six Intel Core i5-9400, 32GB RAM running Ubuntu 20.04.6 and connected by a 1 Gbit switch.
  • ConCap. A Kubernetes cluster (v1.29.0) with 1 control plane and 3 worker nodes. Each node has 16GB RAM, four Intel Xeon E5-2640v4, running Ubuntu 22.04.3. The machines are interconnected by a 10 Gbit switch.
For each experiment, we follow a similar procedure. We first consider the physical network. We instruct one bare-metal host, acting as the “attacker”, to carry out each of the attacks envisioned in the experiment; whereas the other host acts as the “target” and is configured to enable the successful execution of the attack (e.g., if the attack is an SSH bruteforcing, the “target” will be running an SSH server). The corresponding traffic is captured (as PCAP) on the “target” host via tcpdump, and we also generate the corresponding network flows (via CICFlowMeter and Argus). Then, we consider the ConCap environment: we configure ConCap so that the “attacker” and “target” hosts resemble (in terms of computational power, software, and commands executed) of the “bare metal” machines; this ensures that the only “variable” across our experiments is the overarching environment (i.e., ConCap or the physical setup). The configuration files are in our repository repository .
TABLE I: Broad analysis of various network activities. We qualitatively compare the number of packets and NetFlows generated by our bare-metal and ConCap environments. Network traffic on bare-metal servers and ConCap.
Tool Number of Packets CICFlowMeter Flows Argus Flows
Bare-metal ConCap Bare-metal ConCap Bare-metal ConCap
ping 2020 2020 11 11 $10$ $10$
dig $10$ $10$ 55 55 55 55
mysql $37$ $27$ 11 11 11 11
curl/ftp $4910$ $1675$ 22 22 22 22
nmap 55 55 22 22 22 22
nmap (-sV) $128$ $103$ $10$ $10$ $11$ $11$
patator (SSH) 30 96030\,960 31 48531\,485 $680$ $680$ $680$ $680$
patator (FTP) $2093$ $1860$ $70$ $70$ $70$ $70$
slowloris 21 30021\,300 21 30521\,305 $1500$ $1500$ $1500$ $1500$
wfuzz $2310$ $2086$ $15$ $15$ $15$ $15$
V-A Broad Analysis (many different network activities)
We begin our quest to answer RQ1 with a broad, preliminary analysis of quantitative nature. Specifically, we use our environments to reproduce ten network activities of various type, capture the corresponding traffic and generate the NetFlows, and quantitatively compare the overall number of packets and NetFlows generated by both environments.
Network Activities. For a comprehensive assessment, we consider activities of both “benign” and “malicious” nature. Four are clearly malicious: wfuzz (an exemplary HTTP fuzzying attack wfuzz ); slowloris (an exemplary DoS attack slowloris ); as well as two variants of patator (an exemplary bruteforcing attack), one focused on SSH- and another on FTP-bruteforcing. Whereas the other activities are common network-related commands, not necessarily of a malicious nature: ping ping , dig dig , mysql mysql , curl curl over ftp ftp , as well as two variants of nmap nmap . Such a broad set of activities is hence a solid basis to investigate RQ1 at a high level.
Results. We execute each activity with “default” options (we report the exact commands and configurations in the Appendix B-D). For each activity, we capture the corresponding packets and generate the respective NetFlows. We report in Table I the results of this experiment. Specifically, we report the total number of packets, as well as the total number of NetFlows (generated both via CICFlowMeter and Argus) of each activity for the physical and ConCap environments. We can already see that the number of NetFlows is a perfect match, which is a promising result in the context of answering RQ1 positively. Regarding the number of packets, we see that only three activities (ping, nmap, dig) had an equal number of packets. In the next experiment, we will better examine two activities for which we found differences: the SSH variant of patator, and inmapn its service/version probe variant.
V-B In-depth Analysis (multiple variants of two activities)
Our previous experiment (§V-A) revealed that even though ConCap produced nearly-identical traffic (quantitatively-wise) to that of a physical network environment, some network activities (e.g., ssh-patator and nmap) presented some differences. Here, we better examine these phenomena.
V-B1 Methodology
For a deep understanding, we expand our previous assessment and consider 21 variants of the aforementioned activities—up from 2 (i.e., the default configurations). Specifically, for nmap, we consider the 12 possible combinations of the scanning options: -Pn -sS -sV -sT -sU. Whereas for ssh-patator we test 9 combinations by varying: persistent=0/1 -RL=0/1 -T=1/5/10. Note that, for both patator and nmap, our variants also include the default option—which we repeat for completeness. After capturing the PCAP trace and generating the NetFlows, we first examine differences (between ConCap and the physical environment) at the packet level: to this end, we use Wireshark wireshark for qualitative assessments, and then quantitatively inspect the number of packets generated by each variant of our considered activities. Finally, we quantitatively examine differences at the NetFlow level.
V-B2 Packet-level analysis
After qualitatively inspecting the PCAP traces, we found that, across the 12 nmap and the 9 patator variants, the bytes in the individual packets generated by ConCap exactly match those of the physical network—except for expected differences in headers (e.g., MAC and IP addresses, high ports, checksums). We also observed some variations in the window size and maximum segment size which affect the amount of data that can be handled by the receiver: such (minimal) differences are due to the physiological diversity of each network, and their existence is evidence that the packets generated by ConCap can also present a degree of uniqueness which is intrinsic to physical real-world networks.
TABLE II: Network traffic on bare-metal servers and ConCap. For both network activities, the number of NetFlows is identical and the number of network packets has minimal variations (as expected in realistic networks).
Attack Options Number of Packets CICFlowMeter Flows Argus Flows
Bare-metal ConCap Bare-metal ConCap Bare-metal ConCap
-Pn -sS 5 5 2 2 2 2
-Pn -sS -sV 122 103 10 10 11 11
-Pn -sT 6 6 2 2 2 2
-Pn -sT -sV 127 107 10 10 11 11
-Pn -sU 4 4 3 3 4 4
-Pn -sU -sV 4 4 3 3 4 4
-sS 13 13 5 5 6 6
-sS -sV 137 113 13 13 15 15
-sT 14 14 5 5 6 6
-sT -sV 133 115 13 13 15 15
-sU 12 12 6 6 8 8
-sU -sV 12 12 6 6 8 8
-Pn = Treat host as online
-sS / -sT / -sU = TCP SYN, TCP Connect or UDP scan
-sV = Probe open ports to determine service and version info
(a) Nmap port scan
Attack Options Number of Packets CICFlowMeter Flows Argus Flows
Bare-metal ConCap Bare-metal ConCap Bare-metal ConCap
P=0 RL=0 T=1 95 19495\,194 78 30978\,309 $3400$ $3400$ $3400$ $3400$
P=1 RL=0 T=1 33 10333\,103 29 69829\,698 $578$ $578$ $578$ $578$
P=1 RL=1 T=1 38 72638\,726 35 33235\,332 $578$ $578$ $578$ $578$
P=0 RL=0 T=5 95 17095\,170 78 42678\,426 $3400$ $3400$ $3400$ $3400$
P=1 RL=0 T=5 33 55433\,554 30 29330\,293 $595$ $595$ $595$ $595$
P=1 RL=1 T=5 39 06639\,066 35 82335\,823 $595$ $595$ $595$ $595$
P=0 RL=0 T=10 94 94194\,941 78 34778\,347 $3400$ $3400$ $3400$ $3400$
P=1 RL=0 T=10 35 39935\,399 31 50631\,506 $680$ $680$ $680$ $680$
P=1 RL=1 T=10 40 77240\,772 37 38437\,384 $680$ $680$ $680$ $680$
P = Persistent, RL = Rate-Limit, T = Threads
(b) Patator SSH Bruteforce
Next, we focus on the total number of packets exchanged for each activity. We report the results in Tables II(a) (for nmap) and II(b) (for patator). The leftmost column reports the specific options for each attack, whereas the second and third columns report the packets exchanged by the “target” and “attacker” host during the attack for both the physical and ConCap environment (we also report the NetFlows, covered in §V-B3).
  • nmap. For the port scan, there is an almost perfect match. The differences occur only when the attacker probes the service running on the open port (option -sV). This is expected: the -sV option induces the server to provide the index HTML page of the Apache webserver, which is 10,918 bytes. In ConCap, such an exchange requires 2 packets, whereas the same payload requires 8 packets in the physical network—due to small differences in the network conditions such as TCP window scale rfc813 ; rfc7323 .
  • patator. For the SSH brute force, we observe differences of $\approx$$10\%$ in the number of packets. In this case, the difference is due to the TCP/IP stack handling data acknowledgment on the different hardware setups. Additional testing in another replication experiment on a second real network showed similar but different deviations in the number of ACKs. The TPC RFC rfc793 allows for this nondeterministic behavior of “delayed ACKs” which send fewer than one ACK segment per data segment received and is even expected by the official specification rfc813 ; rfc1122 to increase efficiency in the Internet and the hosts. However, as demonstrated by our qualitative analysis, the payload is the same.
In summary, any packet-level differences are due to inherent and unpredictable characteristics of the network, which do not affect the contents of the communication payloads.
V-B3 NetFlow-level analysis
First, we carry out a quantitative comparison focusing on the number of NetFlows exchanged between the “attacker” and “target” host for each activity—reported in Tables II(a) and II(b). Notably, there is always a perfect match: despite differences in packet counts, the number of NetFlows is consistent when the PCAP is processed by the same NetFlow software (CICFlowMeter and Argus follow a different logic to create NetFlows [^3] Bugfix: we encountered an unexpected discrepancy in the NetFlows generated by Argus. We reached out to the developers, and they confirmed that there was a bug in their implementation, and we helped fixed their code. After fixing the code, the number of NetFlows is identical.). Then, we analyze the distributions of NetFlow features (for simplicity, we focus only on CICFlowMeter) across the different attacks. Fig. 4 presents side-by-side comparisons of the mean packet length distribution for all variations of our attacks. We also report the plots for all 30 NetFlow features in Fig. 6 (in the Appendix D). We observe that all features exhibit similar distributions, or any differences can be explained via our packet-level analyses (e.g., more packets sent with empty payload lead to a decrease in the mean packet length).
Figure 4: NetFlow feature distribution of mean packet length for traffic generated by ConCap and a pair of physical hosts. The leftmost plot is for nmap and the rightmost is for patator. Additional plots in Fig. 6.
Answer to RQ1. Our packet- and NetFlow-level analyses revealed that the network traffic generated by ConCap resembles that of a physical network. Deviations are due to expected differences in the network channel, which are impossible to control—but do not affect the payload content.
V-C Determinism of ConCap’s generated traffic
In our previous experiments, we found an intriguing inconsistency. Indeed, by comparing the number of packets generated via the (default) nmap and ssh-patator commands in the first experiment with those of the equivalent veriant (i.e., nmap -Pn -sS -sV and ssh-patator P=1 RL=0 T=10) of the second experiment (cf. Table I with Tables II), we see minor deviations—despite the command and the setup being identical (for both ConCap and the physical network). While it is known that real-world networks exhibit immense variability apruzzese2023sok ; sommer2010outside , and hence such differences may be expected, we do not know if the same also holds for ConCap. Therefore, we ask ourselves our second research question (RQ2):
Research Question #2: To what extent is the traffic generated by ConCap deterministic?
Motivation. There are three reasons why RQ2 is worthy of being investigated. First, to our knowledge, it is still unknown if the traffic generated by containers built on Kubernetes-related frameworks is deterministic; for instance, the authors of DetGen clausen2019traffic tested this hypothesis for Docker, which is different software. Second, because among ConCap’s aims is that of fostering reproducibility (§IV-A), and it is hence crucial to examine the degree of similarity across multiple executions of the same scenario via ConCap. Third, because it enables one to measure the extent to which ConCap can approximate the “unpredictable” behavior of real-world networks.
Setup. To answer RQ2, we created four scenarios (similar to those used in §V-A and §V-B) with increasing complexity in the traffic exchanged between the attacker and target: a simple ping scan (10 ICMP requests), a basic port scan (nmap -sS), a full port scan (nmap -A -T4), and an ssh brute-force attack (patator -P=1 -RL=0 -T=10). We capture the packets and generate the NetFlows for each scenario. We repeat each scenario 100 times. The intention is to study the degree of similarity across all of these repetitions. We also carry out the exact same operations (repeating them 10 times) on the bare-metal servers to compare with a physical setup approximating a real-world network, which should not be deterministic.
Results. Table III reports the results (mean and std) across our trials for both setups, detailing the number of packets, byte count, and the number of NetFlows for CICFlowMeter and Argus. We analyse these results by focusing on ConCap.
  • We see no variation for the ping and basic port scan (nmap -sS) at both the packet- and NetFlow level. This shows that the network connection is reliable: these scenarios involve sending a single request and receiving a single response (or none). Variation would only occur if the network between the attacker and target were unreliable, causing packet duplication, loss, or corruption.
  • In the complex scenarios (nmap -A -T4 and patator -P=1 -RL=0 -T=10), some variation is observed at the packet level, and in the case of the full port scan (nmap -A -T4), even at the NetFlow level. Similar to the realistic traffic assessment, packet-level variation arises from differences in how much data can be transmitted in a single packet and how this data is acknowledged. The flow-level variation in the full Nmap scan is caused by the aggressive mode (-T4) used by Nmap, which can overload the target.
By comparing the results of ConCap with those of the physical network, we see a consistent behavior at the NetFlow-level. The differences are only for the Full Port Scan and amount to less than 0.1%, which are not significant. In contrast, more pronounced differences exist for the Full Port Scan and SSH Bruteforce attacks from a packet-level viewpoint. In these cases, ConCap generates an average of $\approx$9% less packets and $\approx$9% less bytes than the physical setup.
Answer to RQ2. Traffic generated by ConCap is deterministic content-wise, but the nondeterministic nature of networking results in small variations in packets and bytes. Variations are due to how data is acknowledged (e.g., using a separate TCP packet) and should not impact the outcome of a scientific experiment. ConCap does not just simulate network traffic, it generates it such that it resembles physical networks—whose real-world behavior is unpredictable sommer2010outside .
TABLE III: RQ evaluation. We repeat the scenarios with ConCap 100 times.
Environment Attack Packets Number of Flows
Count Sum of Bytes CICFlowMeter Argus
Bare-Metal Ping scan $20\pm 0.$ $1960\pm 0.$ $1\pm 0.$ $10\pm 0.$
Basic Port Scan $13\pm 0.$ $736\pm 0.$ $5\pm 0.$ $6\pm 0.$
Full Port Scan $2751\pm 88.$ $271\,551\pm 6235.$ $1091\pm 43.$ $1093\pm 43.$
SSH Bruteforce $30\,935\pm 37.$ $5\,060\,797\pm 2417.$ $680\pm 0.$ $679\pm 0.$
ConCap Ping scan $20\pm 0.$ $1960\pm 0.$ $1\pm 0.$ $10\pm 0.$
Basic Port Scan $13\pm 0.$ $694\pm 0.$ $5\pm 0.$ $6\pm 0.$
Full Port Scan $2504\pm 6.$ $239\,401\pm 4631.$ $1098\pm 1.$ $1092\pm 1.$
SSH Bruteforce $26\,960\pm 354.$ $4\,759\,703\pm 23\,353.$ $680\pm 0.$ $679\pm 0.$
VI Applications: ConCap for Research
Here, we show the utility of ConCap via proof-of-concept experiments involving noteworthy practical applications—especially from an ML viewpoint. First, we show that ConCap can be used to replicate prior research (§VI-A). Then, we show that ConCap enables creation of ML-based NIDS from scratch (§VI-B). Finally, we show how to use ConCap for assessments of existing NIDS against unseen attacks (§VI-C).
Disclaimer: the following “proof of concept” experiments are just meant to demonstrate some applications of ConCap. We do not claim generalizable results, nor seek to cover all possible cases.
VI-A Replicability of Prior Research
Recall that our previous experiments (§V) revealed some small differences (at least at the packet level) in the data generated by ConCap w.r.t. of a physical network. Hence, we ask ourselves: are such differences “significant” from the viewpoint of ML-based network-intrusion detection?
Goal. We explore this question by attempting to reproduce the results of prior work focusing on ML-based NIDS. Specifically, we consider the experiments carried out by Engelen et al. engelen2021troubleshooting and Liu et al. liu2022error , which involved assessment of supervised ML-based classifiers analysing network flows created via a (fixed) version of CICFlowMeter (the same one we used for ConCap). Specifically, the NetFlows pertain to the well-known CICIDS17 and CICIDS18 benchmark datasets sharafaldin2018toward ; such NetFlows have been rigorously labeled by the authors of engelen2021troubleshooting ; liu2022error . Both CICIDS17 and CICIDS18 include, among others, network-traffic data generated with ssh-patator (which we used in §V). Our intention is to train various ML models on the NetFlows of ssh-patator included in CICIDS17 and CICIDS18, and then testing such classifiers on the NetFlows of ssh-patator generated via ConCap, and then comparing the results: our expectation is that the classifiers trained on either “version” (i.e., CICIDS17/18, or ConCap) of ssh-patator shall always be able to detect the other version—thereby demonstrating that the data generated by ConCap is functionally equivalent (ML-wise) to that of prior research.
Setup. We follow the instructions provided in liu2022error ; engelen2021troubleshooting , and retrieve the (fixed) version of CICIDS17 and CICIDS18 (more details in Appendix C-B). Then, for each dataset, we partition the NetFlows according to their classes. For this experiment, we are only interested in the ssh-patator class (which, in both CICIDS17 and CICIDS18, was launched with the default options), which we denote as $\mathcal{P}$; and the benign class, which we denote as $\mathcal{B}$. Then, we take the NetFlows generated by ConCap with the default ssh-patator (see §V-A), which we denote as $\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$. For a broad assessment, we consider five different ML-based classifiers: Random Forest (RF), Decision Tree (DT), XGBoost (XGB), Support Vector Machine (SVM), as well as a Deep Neural Network (DNN); these classifiers have been tested in liu2022error ; engelen2021troubleshooting , typically obtaining near-perfect performance. Then, for each dataset, we train each classifier on 80% of $\mathcal{P}$ and 80% of $\mathcal{B}$, and test it on the remaining 20% of $\mathcal{B}$ (for the false positive rate, fprfpr), the remaining 20% of $\mathcal{P}$, and all of $\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$ (computing the true-positive rate, tprtpr for both sets). We then repeat this process, but by switching $\mathcal{P}$ with $\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$. We repeat this procedure 5 times for statistical robustness. Note that such a workflow complies with the best practices of ML-based assessments in cybersecurity and ML-NIDS arp2022and ; apruzzese2023sok . The experimental source-code is in our repository repository .
TABLE IV: Reproducibility of prior work. We show ConCap-generated data is functionally equivalent to that of existing benchmark datasets.
Train Set     $\mathcal{B}$$\mathcal{P}$     $\mathcal{B}$$\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$
Test Set     $\mathcal{B}$ $\mathcal{P}$ $\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$    $\mathcal{B}$ $\mathcal{P}$ $\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$
CICIDS17 DT     <<0.001 0.997 0.917    <<0.001 0.980 1.000
RF     0.000 0.998 1.000    0.000 0.986 1.000
XGB     <<0.001 0.998 0.869    <<0.001 1.000 1.000
SVM     <<0.001 0.993 0.907    <<0.001 0.993 1.000
DNN     0.000 0.998 0.920    <<0.001 0.996 1.000
CICIDS18 DT     0.000 1.000 0.859    <<0.001 1.000 1.000
RF     0.000 1.000 0.859    <<0.001 1.000 1.000
XGB     0.000 1.000 0.859    <<0.001 0.984 1.000
SVM     0.000 1.000 0.907    <<0.001 1.000 1.000
DNN     0.000 1.000 0.907    <<0.001 1.000 1.000
Results. We report the results of our tests in Table IV. The table shows the fprfpr, which is always less than 0.001 (a result consistent with prior work liu2022error ; engelen2021troubleshooting ), even when the classifiers were trained with ConCap’s generated data. Then, looking at the tprtpr, we can see that the results on $\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$ align with those on $\mathcal{P}$, both on CICIDS17 and CICIDS18, and irrespective of whether the training was done on 80% of $\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$ or $\mathcal{P}$. Indeed, despite not being “perfectly” aligned [^4] We conjecture that such minor difference may stem—besides from unpredictable network effects—from potential labeling inaccuracies in the (fixed) CICIDS17 or CICIDS18, since these tasks have been done manually (whereas ConCap does so automatically); they can also be due to different machines (the creators of CICIDS17/18 do not provide low-level hardware details sharafaldin2018toward ). the tprtpr always shows that these attacks can be detected, since the tprtpr is always above 0.85. Importantly, this result is shared across all models/classifiers. Thus, we can say that, from an ML viewpoint, and at least according to these experiments, the (labeled) NetFlows generated via ConCap can be used to replicate prior research and derive the same conclusions.
VI-B Development of ML-based NIDS for the Real World
We consider the case in which a researcher may want to develop an ML-based NIDS from scratch. We assume that the researcher has access to a real-world network and that the hosts of such a network are not compromised. In this context, we wonder: can ConCap be used to develop an ML-NIDS that detects malicious traffic generated by real-world networks?
Goal. This experiment is conceptually similar to the one discussed in the previous subsection (§VI-A). At a high level, we will use ConCap to generate malicious network activities, and do the same on a real-world setup; and then see if using the malicious NetFlows generated via ConCap to train an ML-based classifier yields a model that can detect the real-world version of the malicious network activities—and vice versa. The difference, however, is that we are not relying on benchmark datasets here: this experiment is carried out on a real-world “smart home” network. [^5] Yes, we captured traffic from a network regularly used by real people, and infected hosts deployed in such a network. We obtained permissions by the users/residents. We discuss ethical considerations at the end of the paper. Our intention is twofold: first, verify if the results we achieved in our previous experiment also hold here—thereby reinforcing the claim that the data generated by ConCap is unlikely to alter the conclusions of a ML-based experiment; second, examine how well the model trained with ConCap’s NetFlows can detect their real-world variant (which technically represents the “true attack” that the ML-NIDS should protect against).
Setup. For this experiment, we consider a real, physical network encompassing $\approx$50 active hosts (note: this network is different from the one of the “bare-metal” servers used in §V). These hosts include a mix of IoT devices, gaming consoles, laptops, as well as smartphones. We captured $\approx$20GB of network traffic in this network, which we verified to be clean of malicious activities. To capture the malicious activities, we consider two hosts (i.e., two laptops) deployed in this network, one acting as the “target” and the other as the “attacker”. To align this experiment with the one in §VI-A, we consider the ssh-patator as our attack. Hence, we deploy an SSH server on the “target” host, and use the “attacker” host to launch ssh-patator (with its default options) against the SSH server. We captured the corresponding PCAP trace (on the “target” host), extracted the NetFlows (via CICFLowMeter) and manually labeled them. We then follow the same workflow as in §VI-A for the ML-based experiments, using the same notation. These experiments are also available in our repository repository and more details on the overarching real-world network and packet capture are provided in the Appendix C-A.
Results. The results of this experiment are in Table V. We can see a similar trend to that shown in the previous experiment (cf. Table IV with Table V). In particular, the models trained with ConCap-generated data ($\mathcal{B}$$\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$ columns) have a near-zero fprfpr and a near-perfect tprtpr, both on the malicious NetFlows generated via ConCap (i.e., $\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$), which is expected; and on those generated by the physical hosts (i.e., $\mathcal{P}$). Moreover, such a good performance also encompasses the case wherein the models are trained only with real-world data ($\mathcal{B}$$\mathcal{P}$ columns). Altogether, these finding further confirm that ConCap can be used to carry out ML- and NetFlow-based experiments (thereby reinforcing our conclusions related to §VI-A); and also show that ConCap can be used to develop ML-based NIDS meant to be deployed in the real world.
TABLE V: Real-world equivalence. Network traffic generated with ConCap is compatible with a real-world ”smart home” network.
Train Set     $\mathcal{B}$$\mathcal{P}$     $\mathcal{B}$$\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$
Test Set     $\mathcal{B}$ $\mathcal{P}$ $\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$    $\mathcal{B}$ $\mathcal{P}$ $\mathcal{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$
DT     <<0.001 1.000 0.899    <<0.001 >>0.999 >>0.999
RF     0.000 1.000 0.899    0.000 1.000 1.000
XGB     0.000 >>0.999 0.889    <<0.001 1.000 1.000
SVM     0.000 1.000 1.000    <<0.001 1.000 1.000
DNN     <<0.001 1.000 1.000    <<0.001 1.000 1.000
VI-C Security Assessment of Existing (ML-based) NIDS
Our previous experiments showed that ConCap’s generated data can be used to develop ML-based NIDS, showing that, by training classifiers on some malicious data of a “known” attack, it is possible to detect future instances of the same attack. Here, we show how ConCap can be used to produce new knowledge. Specifically, we ask ourselves: how can ConCap be used to test existing NIDS against “unseen” attacks?
Threat Model. We consider a defender that uses an ML-NIDS developed by using either the CICIDS17 or CICIDS18 dataset (in their fixed version engelen2021troubleshooting ; liu2022error ); such an assumption serves for broad coverage of diverse use cases, since these datasets are well-known “benchmarks”, encompass a variety of attacks, and are used even in recent and top-tier publications (see, e.g., flood2024bad ). Differently from our previous experiments (in §VI-A) here we use all of the malicious data (not just that of ssh-patator) in these datasets to develop our models, leading to much larger, and therefore more effective, detectors. The attacker seeks to evade the detection by using recent exploits, whose corresponding network traffic has not been included in the training dataset of the detector used by the defender. Without loss of generality, we consider three real and recently published exploits, taken from the common vulnerabilities and exposures (CVE) database: CVE-2024-47177 cve1 (related to OpenPrinting Cups), CVE-2024-36401 cve2 (related to GeoServer), CVE-2024-2961 cve3 (related to GNU C Library). The attacker hence exploits these CVEs, and hopes that such malicious activities are not detected by the defender.
Approach. We train five ML-based classifiers (RF, DT, SVM, DNN, XGB) on the whole CICIDS17 and CICIDS18 datasets; we confirm that such ML-NIDS obtain the same performance as prior work engelen2021troubleshooting ; liu2022error . Then, we use ConCap to generate network-traffic data of the three considered CVEs. We setup ConCap accordingly, so that the target host can be “exploited” via the commands included in the CVE and executed on the attacker host. To expand our coverage (and also to show the capabilities of ConCap), we perform the captures by varying the overarching network conditions, specifying different values for delay, loss, corrupt, and duplicate, totaling over 320 combinations. Indeed, from a practical viewpoint, ConCap makes such repetitions trivial to carry out, and it is sensible to assume that an attacker may exploit any given CVE randomly, i.e., when the network conditions may be subject to bursts or noise. Overall, we capture 350k packets, corresponding to 4800, 640, and 1280 NetFlows for CVE-2024-47177, CVE-2024-36401, CVE-2024-2961, respectively. We then test our 10 models (5 per dataset) on these NetFlows.
Results. We report the results in Table VI (Appendix D). Our models are unable to reliably recognize the CVE NetFlows as malicious (the tprtpr is always below 0.6, and below 0.25 in most cases). Such an insight can be used as a signal that the ML-NIDS must be retrained—which is a trivial task since it is simply necessary to enhance the training dataset with the data generated by ConCap and re-train the ML models by, e.g., leveraging the well-known adversarial training technique apruzzese2023sok . (In Appendix D-B, we also demonstrate a full-fledged multi-step and multi-host attack by using ConCap.)
VII Discussion and Critical Analyses
We evaluate the computing requirements needed to run ConCap (§VII-A), discuss the limitations of our research (§VII-B), and draw lessons learned from our experiments (§VII-C).
VII-A Computing Requirements and Operational Runtime
We show that ConCap can seamlessly run also on commodity hardware (e.g., laptops) by measuring its operational runtime.
Approach. We test ConCap by performing various experiments on two setups: a multi-node Kubernetes cluster (the same used in §V), and a laptop with an Intel i7-1265CPU and 23GB RAM (more hardware details in the Appendix D-A). Specifically, we consider the 12 combinations of nmap (also used in §V-B). Note that what we are interested in is the background utilization of resources (CPU and RAM) as well as the overall initialization time, i.e., the time required before the “attacker” host begins issuing its commands. Anything beyond this step does not depend on ConCap (e.g., if the attacker wants to DDoS the target, or if the scenario requires setting up a target running resource-intensive services, then such requirements are unrelated to ConCap’s resource utilization).
Results. We report the results in Table VIII (in the Appendix D-A, which also provides more details). On average, less than 3 seconds elapse (on both setups) before the nmap command is launched. We also see that, on both setups, the containers running nmap requires less than 2MiB of RAM (across all of our experiments); whereas ConCap’s memory footprint is less than 40MB. In terms of CPU, the utilization varies (which is expected, since it depends also on the dynamic allocation of the machine) but, on average, the max-CPU used was 10% for the cluster and 6% for the laptop. We can hence conclude that ConCap is also intrinsically lightweight.
VII-B Scope and Limitations
We advance the state of the art by providing a tool, ConCap, to foster future research on NID. Our contribution are aimed at research: real-world deployments are outside our scope (as we clearly remarked in §II). Although we did rely on real-world assessments, such assessments primarily served to prove the utility of ConCap for this specific purpose.
As for limitations, we acknowledged ConCap is not fully deterministic: while this property can be a strength (e.g., by running the same scenario multiple times, it is possible to collect more “realistic” data) it can also be a weakness (e.g., for reproducibility of some experiments expecting to achieve perfect matches). However, such a limitation is expected due to the non-deterministic nature of modern networks/protocols (see §V-B2). Second, ConCap’s labeling is addressed to network flows, and cannot hence be used for labeling of other datatypes which can be used for NIDS purposes (e.g., venturi2023arganids ; yet, as we argued (in §IV-A) the majority of papers on NIDS rely on network flows for their analyses. Third, and related to the experiments carried out in this work, we believe the answer to our major research questions (in §V) to be correct and also generalizable (for instance, even the authors of DetGen clausen2019traffic obtained similar results to ours, albeit with different software). In contrast, we acknowledge that the experiments in §VI cannot be used to cover all possible use-cases: this is why we emphasized that the experiments in that section serve as a “proof of concept” to show potential applications of ConCap.
Finally, we emphasize that ConCap can only generate data according to the specifications provided by the developer.
VII-C Lessons Learned
Let us summarize the major takeaways of our work.
First, ConCap automatically generates network-traffic data conforming to “any” attack. Such data resembles that of a physical network, including its non-deterministic nature (§V). We manually verified all data labeled by ConCap, and the labeling was always correct: the NetFlows always pertained to the malicious commands specified in the configuration file. Such a verification is confirmed by our experiment showing that such data is functionally equivalent to that of benchmarks/real-world networks (§VI-A and §VI-B).
Given the above, ConCap can be used in a variety of ways (some of which shown in §VI). For instance, through ConCap, it is possible to produce new knowledge (or test new hypotheses) without the need (and risk) to carry out experiments entailing malicious activities in real-world networks. Crucially, and as a byproduct of the experiments in §VI-C, ConCap enables early-assessments of existing NIDS (not-necessarily reliant on ML methods) against recent threats. For instance, through ConCap, the owners of an NIDS can safely test their systems by (i) taking any new CVE and (ii) running it on ConCap, and then (iii) either replay the PCAP trace or submit the NetFlows to the NIDS. Depending on the outcome, it is then possible to enhance the NIDS by, e.g., using adversarial training apruzzese2023sok with ConCap’s generated data.
Finally, through ConCap, future research does not need to “share data” (which can require abundant storage space), rather “sharing the scenario configuration file” (typically less than 1KB in size) and the experimental details (e.g., the CVE used) is sufficient to reproduce prior results.
Q&A: We clarify some remarks (e.g., fitness to SaTML and realism of ConCap’s data) made by reviewers in the Appendix A-C
VIII Related Work
We are not aware of any open-source tool that fulfills the same goals as ConCap. Here, we first summarize related literature (§VIII-A), then compare ConCap with the two closest works we could find (§VIII-B), and finally suggest avenues in which ConCap can benefit future related research (§VIII-C).
VIII-A Literature Summary
Some orthogonal works propose testbeds that do not provide the functionalities of ConCap: e.g., Gotham saez2023gotham cannot ensure fine-grained labeling of malicious datapoints, whereas I2DT cordero2021generating can only inject packets in a PCAP.
Closely related works are DetGen clausen2019traffic and SOCBED uetz2021reproducible ; we tried to find more works on “generation of realistic network traffic” by looking at the accepted papers in various top-tier conferences (NDSS, IEEE S&P, USENIX Sec, ACM CCS) since 2019 (similarly to arp2022and ; apruzzese2023real ). We found that the only related work was netUnicorn beltiukov2023search . Then, we expanded our search via snowballing wohlin2014guidelines and (recursively) looked for all peer-reviewed papers cited by clausen2019traffic ; beltiukov2023search , and identified four related works beltiukov2023pinot ; khan2024harnessing ; buhler2022generating ; zhoutrafficformer ; guthula2023netfound . Let us position ConCap within these related works.
PINOT beltiukov2023pinot and netUnicorn beltiukov2023search , can generate benign traffic but raise security risks for generating malicious traffic (see §II-B). To generate benign traffic at scale, netMosaic khan2024harnessing ; buhler2022generating harnesses public code repositories to automatically capture network traffic for a wide range of applications, but does not allow fine-grained labeling of malicious datapoints. Finally, Zhou et al. zhoutrafficformer propose to use foundational models to artificially “augment” any given network-traffic dataset; however, it is unclear if such methods can produce data resembling a physical setup (there is no real-world assessment in zhoutrafficformer ).
VIII-B Low-level Comparisons
We compare ConCap with DetGen detgenGithub and SOCBED socbedGithub .
ConCap vs DetGen. First, DetGen clausen2019traffic does not allow parallelism by design. Second, DetGen does not generate and label the NetFlows of any given experiment, thereby forcing the developer to do so manually—which is error-prone engelen2021troubleshooting . Finally, DetGen does not allow the same degree of flexibility provided by ConCap. To provide evidence of this claim, we looked at the public repository of DetGen (available at detgenGithub ), inspected its source code, and observed the following: (i) we did not find any way to set the available CPU/RAM of the attacker/target; (ii) a predefined timeout is required for every scenario, potentially stopping the scenario before successful completion; (iii) comments in the code state that the attack sometimes starts before traffic is captured, highlighting the lack of fine-grained control over the scenario execution by DetGen. Although we tried to compare the traffic generated by both approaches, we were unable to run any of their example scenarios (due to deprecated software/runtime errors that we could not troubleshoot with the provided documentation).
ConCap vs SOCBED. First, SOCBED primarily targets log-data collection: even though it can create PCAPs, it does not (a) extract NetFlows by default, or (b) label them. The latter is crucial: manual post-hoc labeling of attack traffic is unreliable due to the background noise generated by the concurrent simulation of benign and malicious behaviors. Second, SOCBED relies on a fixed network topology designed for long-running experiments running virtual machines on a single host. Their sample scenario requires around one hour to complete, including a 15-minute setup time. In contrast, ConCap supports fast (startup time of seconds), parallel execution of isolated attack scenarios across multiple hosts. Finally, we found no example to configure network characteristics, such as per-host bandwidth or delay, in SOCBED.
Replication: We show in the Appendix B-B and B-C how to replicate one of DetGen’s and SOCBED scenarios with ConCap.
VIII-C Future Work
Future research can use ConCap to investigate open problems in NIDS ceschin2024machine ; cina2024machine , such as: robustness to concept drift andresini2021insomnia ; wang2022enidrift , explainability bhusal2023sok ; mink2023everybody ; wei2023xnids , false alarms alahmadi202299 ; van2022deepcase ; vermeer2023alert ; fu2023point , or development of novel detection techniques (not necessarily relying on NetFlows, such as mirsky2018kitsune ; sharif2024drsec ; rehman2024flash ; yang2023prographer ). ConCap can technically also be used for generation of benign labeled traffic (useful for, e.g., traffic classification azab2024network ; cebere2024understanding ; siracusano2022re ; rimmer2022trace ; schafer2022accurate ). To further demonstrate the broad applicability of ConCap, we use its data to test other detection approaches wei2023xnids ; luay2025multimodal , based on deep learning or transformers, in Appendix D-C.
IX Conclusions
Our paper is a stepping stone for future research in NID.
With ConCap, we enable researchers—especially those without access to real-world networks—–to conduct realistic security assessments in networking contexts. ConCap automatic (and correct) labeling capabilities remove the burden of carrying out manual annotation. Moreover, ConCap open-source imprint fosters reproducibility and transparency.
Acknowledgments
The authors would like to thank the anonymous SaTML’26 reviewers for the feedback received, which substantially improved our work. Part of this research was funded by the Hilti Corporation. This research was also partially supported by the BOSA-AIDE project funded by the Belgian SPF BOSA with reference number 06.40.32.33.00.10. This work originated from a research stay at the University of Liechtenstein and was supported by the Research Foundation – Flanders (FWO) under grant number V450224N.
Appendix A Ethics, Open Science, and Clarifications
Following the best scientific practices, and to also comply with the CfP, in this appendix we: provide some ethical comments (Appendix A-A); describe the technical content we will publicly release (Appendix A-B); and elaborate our response to some valid remarks that we received in previous versions of this paper (Appendix A-C), while also disclosing our “responsible usage” of compute.
A-A Ethical Remarks
We make the following ethical considerations.
Prevention of harm. For our research, we carried out experiments entailing “network attacks” (i.e., ssh bruteforcing patator ) on a real-world network. The authors received permission to do so, and no software/hardware was damaged as a result of these operations.
Data Confidentiality. For our research, we have collected data from two physical networks: that of the bare-metal setup, and the real-world one encompassing $\sim$50 hosts. We will release the full data (PCAP and NetFlows) for the bare-metal setup. For the other network, we will only provide the (labeled) NetFlows, which do not leak private data (we will anonymise the IP addresses and ports). We received permission by the owners to share this data publicly.
Respectfulness. We emphasize that our intention is not to “point the finger” against any prior work. We found a bug in the current implementation of DetGen detgenGithub that prevented us from proceeding in our comparison (see §VIII). We contacted the developers of DetGen, informing them of the issue. They responded and confirmed that the current code is outdated and does not function properly. We are currently corresponding with them, offering our help to fix the problem.
A-B Data Availability
All necessary artifacts to replicate the results in the paper will be released together with ConCap allowing researchers to create their own traffic generation experiments. Specifically, all our resources are available at: https://github.com/idlab-discover/ConCap repository (an anonymised version of this repository was also provided for peer review). The repository contains the following resources:
  • The source code of ConCap (which is based on open-source libraries)
  • The containers used for our experiments.
  • The PCAP and (labeled) NetFlows generated by ConCap that we used for our experiments.
  • The PCAP trace and NetFlows of the bare-metal setup.
  • The labeled NetFlows of our real-world network of 50 hosts; (we cannot release the PCAP for privacy).
  • The notebooks of our experiments (for reproducibility).
  • The additional data (PCAP and labeled NetFlows) used to showcase that ConCap can be used to generate “new” data conforming to recent attacks.
We have the permission to share all of the above.
A-C Clarifications on our Research
We provide additional clarifications on four aspects: the labeling accuracy of ConCap, the realism of ConCap’s generated data, the necessity to provide ad-hoc configuration files to run ConCap, the resources used to carry out our experimental evaluation, and fitness of our work to the SaTML community. To facilitate understanding, we organize this appendix in a “Question-and-Answer” format.
How does ConCap ensure that the NetFlows are correctly labeled? This is clarified in Section III-B. The major difficulty in labeling network data (and, in our case, NetFlows) is that it is difficult precisely distinguish malicious datapoints among the myriad of activities carried out by modern machines. For instance, the labeling of most popular benchmarks is done by (i) generating some network traffic in synthetic environments, (ii) creating the corresponding NetFlows, and (iii) apply coarse labeling strategies—such as “assign the label \$maliciousLabel to all NetFlows generated by \$IPaddress between \$startTime and \$endTime using ports \$dPort and \$sPort” (see flood2024bad ). Such strategies may inevitably lead to labeling as malicious also benign/background traffic, and they may also underestimate the actual malicious traffic (e.g., there may be malicious activities related to a given malware that entail ports different from \$dPort or \$sPort”). In contrast, the labeling applied by ConCap does not have such a problem by design: thanks to its isolated environment and to the fact that the “attacker” host runs only the (supposedly malicious) commands specified in the configuration file, it is guaranteed that the packets generated by the attacker host will include only the network activities that pertain to the provided command(s). In our experiments, we have verified that this holds true (see §VII-C); moreover, the fact that the ConCap’s generated-and-labeled data yielded ML-NIDS that have the same performance as in prior work (see experiment in §VI-A) further confirms that ConCap provides accurate labeling—by design and without requiring human intervention.
ConCap generates network traffic in a synthetic environment. How can such traffic be realistic? From a technical viewpoint, network packets have no notion of “real-world” or “synthetic” environment. Therefore, the payload (i.e., the actual data that flows over a network and that is exchanged between two endpoints) should be identical whether the traffic capture occurs in a “synthetic” or “physical” testbed. However, networks—irrespective of “where” they are—are characterized by having immense variability sommer2010outside . For instance, two organizations having the exact same hosts which carry out the exact same activities can have different network traffic data because, e.g., they may have different bandwidth. Thanks to ConCap’s configuration file, it is possible to specify parameters of the network channel that can lead to a better approximation of the real-world network behavior. Note, however, that from the viewpoint of the malicious payload, we have empirically verified (in §V) that there are no differences.
ConCap requires manual effort to define the configuration files. Wouldn’t this offset the benefit of providing automatic traffic generation and labeling? We respectfully disagree. The analysis done by Flood et al. flood2024bad (and, previously, also by Liu et al. liu2022error and by Engelen et al. engelen2021troubleshooting ) show that manual labeling requires extensive effort. In contrast, with ConCap, it is just necessary to, e.g., read the documentation of a given CVE and setup the hosts accordingly to produce a “configuration file” that can be used as a blueprint to carry out essentially endless variants of a given malicious activity. For instance, for the attacks we captured in §VI-C (and also in Appendix D-B), we only took $\approx$2 hours (we did them in the timespan between a “rebuttal phase” of a security conference, and we can provide evidence of this). Such effort is objectively smaller than that required to, e.g., set up a full-fledged network environment (even via virtual machines), capture the traffic, generate the NetFlows, and precisely label such NetFlows.
What is the memory footprint of your research? Our experiments did not use a lot of compute (we report the training times of our models in Appendix D) and ConCap requires a negligible amount of computing resources to run (see §VII-A). We do not believe more experiments are necessary to prove any of our major claims: doing so (e.g., to show how LLMs can benefit from ConCap’s data) would be a waste of resources for the purpose of this specific paper.
Can you clarify how ConCap represents a significant contribution to the SaTML research community? First, we acknowledge that our paper’s primary contribution, ConCap, is primarily an engineering effort. However, as shown by the EuroS&P’24 Best Paper Award flood2024bad (and also by other recent works, e.g., ACSAC’25 wangr2025rr ), the ML-NIDS research community is in desperate need of “better datasets”. Our contribution, ConCap, is our response to such a need, given that it solves the labeling issue while generating valid data for ML-NIDS research experiments. Moreover, we stress that the development of ConCap goes beyond simply deploying containers. Configuring and executing reproducible, labeled network experiments end-to-end—especially in the context of security research—is not supported by Kubernetes out of the box. For instance, Kubernetes has plugins for networking of containers, but none provides support for traffic shaping (e.g., controlling bandwidth or latency). Moreover, Kubernetes does not natively support packet capture, NetFlow generation, or any form of labeling—the latter being the major problem affecting existing NIDS datasets flood2024bad . While some of these steps may seem trivial in isolation, executing them reliably and reproducibly across scenarios is a known challenge. To sum up, ConCap is a tool to “unlock” new research to address the many open problems in the ML-NIDS domain: without “trustworthy” (which we intend as “correctly-labeled”) data, it is difficult to provide convincing solutions. ConCap therefore enables trustworthy data curation in a way that is safe to practically use (since the capture occurs in a safe setup), both of which being themes that align with SaTML’s vision.
Appendix B ConCap Configuration
We provide information for practical use of ConCap. First, the full configuration files for a ConCap scenario and NetFlow extractor are given with all the possible configuration options. Then, a step-by-step guide is provided to replicate a scenario from DetGen clausen2019traffic .
Listing 1: A ConCap scenario configuration file describing a full port scan via nmap against an Apache webserver.
attacker:
name: nmap
image: instrumentisto/nmap:latest
atkCommand: nmap \$TARGET_IP -A -T4
atkTime: 30s
cpuRequest: 100m
memRequest: 100Mi
cpuLimit: 500m
memLimit: 500Mi
target:
name: httpd
image: httpd:latest
filter: host \$ATTACKER_IP and host \$TARGET_IP and not arp
cpuRequest: 100m
memRequest: 100Mi
cpuLimit: 500m
memLimit: 500Mi
network:
bandwidth: 100mbit
queueSize: 100ms
limit: 10000
delay: 0ms
jitter: 0ms
distribution: normal
loss: 0%
corrupt: 0%
duplicate: 0%
seed: 0
labels:
label: 1
category: port-scan
subcategory: nmap
scenario: nmap_A_T4
B-A NetFlow Exporter Configuration
Automatic NetFlow extraction in ConCap is set-up by processing configuration files. A flow extractor is created for each file, which exports NetFlows from the scenario’s network traces. An example configuration file in Listing 2 for Argus has 3 configuration options: a “name”, “containerImage”, and “command”. The name and container image are used to deploy the NetFlow exporter container. The command is responsible for processing the network capture file and outputting the extracted network flows as a CSV file.
Listing 2: Argus Processing Definition for ConCap
name: argus
containerImage: ghcr.io/idlab-discover/concap/argus:latest
command: argus -r \$INPUT_FILE -S 60s -w - | ra -r - -c, > \$OUTPUT_FILE
B-B Configuration of a DetGen Scenario in ConCap
To demonstrate the flexibility of ConCap, we have created a step-by-step guide to implement one of the predefined scenarios from DetGen clausen2019traffic . We selected the well-documented “capture-020-nginx” scenario, which uses “siege”, an HTTP load testing and benchmarking tool, to target the “nginx” HTTP server and reverse proxy. In this guide, we show how to replicate the “capture-020-nginx” scenario in ConCap using processing pods and scenario definitions.
NetFlow Configuration DetGen does not support automated NetFlow generation, thus we skip the processing pods.
Scenario Configuration The ConCap scenario is detailed in Listing 3. Here, the attacker is configured to use siege for 10 seconds with 10 simulated users, each making 1,000 requests to the reverse proxy’s index page. The IP address of the target is assigned through environment variable expansion. Since there is no official Docker image for siege, we could have opted for one of the many community-built images. However, to demonstrate the simplicity of creating a custom image, we provide a minimal Dockerfile in Listing 4, which we then push to our GitHub Container Registry. For the target, we use the official nginx DockerHub image. To execute the scenario twice, we duplicate the scenario definition file.
Listing 3: The replicated capture-020-nginx scenario in ConCap.
attacker:
name: siege
image: ghcr.io/idlab-discover/concap/siege:ubuntu18
atkCommand: siege -c 10 -r 1000 -v http://\$TARGET_IP
atkTime: 10s
target:
name: nginx
image: nginx:1.13.8-alpine
Listing 4: A minimal container image for running siege.
FROM ubuntu:18.04
ENV DEBIAN_FRONTEND noninteractive
RUN apt-get update && \
apt-get -y install siege && \
apt-get clean && \
rm -rf /var/lib/apt/lists
ENTRYPOINT ["siege"]
Listing 5: A ConCap multi-target scenario file defining a multi-step attack chain against three targets. The scenario combines global with target-specific network and labeling configuration.
type: multi-target
name: multi-step-attack
attacker:
name: advanced-attacker
image: ghcr.io/idlab-discover/concap/advanced-attacker:1.0.0
atkCommand: ./multi-step-attack.sh \$TARGET_IP_0 \$TARGET_IP_1 \$TARGET_IP_2
cpuRequest: 100m
memRequest: 250Mi
targets:
- name: openssh
image: ghcr.io/idlab-discover/concap/openssh-server:password-24.04
cpuRequest: 200m
memRequest: 200Mi
labels:
step: bruteforce-ssh
network:
bandwidth: 10Mbit
queueSize: 100ms
delay: 100ms
- name: db
image: ghcr.io/idlab-discover/concap/mysql:1.0.0
cpuRequest: 200m
memRequest: 200Mi
labels:
step: exfiltration
network:
bandwidth: 1Gbit
queueSize: 100ms
delay: 1ms
- name: libssh
image: vulhub/libssh:0.8.1
cpuRequest: 200m
memRequest: 200Mi
labels:
step: exploit-cve
network:
bandwidth: 100Mbit
queueSize: 100ms
delay: 5ms
labels:
label: 1
category: advanced-lateral
B-C Configuration of a SOCBED Scenario in ConCap
We further show the coverage of ConCap by providing the configuration files that enable reproduction of the experimental setup of SOCBED uetz2021reproducible .
To this end, we report in Listing 6 the configuration that reproduces a portscan, i.e., the first step of the attack chain in SOCBED. We also provide the corresponding PCAP and (labeled) NetFlow data in our repository (note: this data is not provided by SOCBED). Furthermore, in our repository repository , we report additional configuration files (and corresponding data) for other scenarios in SOCBED.
We observe that creating all these configuration files required us less than one day of work: we merely investigated the public repository of SOCBED socbedGithub to infer the low-level network and command details, required to devise our own configuration files.
Listing 6: A configuration of ConCap to support one of the scenarios envisioned in SOCBED which includes a portscan.
type: multi-target
name: socbed-example
attacker:
name: attacker
image: instrumentisto/nmap
atkCommand: nmap \$TARGET_IP_1 \$TARGET_IP_2 \$TARGET_IP_3 \$TARGET_IP_4 \$TARGET_IP_5 -n disable-arp-ping -sU -sV
cpuRequest: 200m
memRequest: 200Mi
targets:
- name: dmz-server
image: vulnerables/web-dvwa
cpuRequest: 200m
memRequest: 200Mi
labels:
server: dmz-server
- name: log-server
image: kibana:9.2.1
cpuRequest: 200m
memRequest: 200Mi
labels:
server: log-server
- name: internal-server
image: dperson/samba
cpuRequest: 200m
memRequest: 200Mi
labels:
server: internal-server
- name: client-1
image: dockurr/windows
cpuRequest: 200m
memRequest: 200Mi
labels:
server: client-1”
- name: client-2
image: dockurr/windows
cpuRequest: 200m
memRequest: 200Mi
labels:
server: client-2”
network:
bandwidth: 100Mbit
queueSize: 100ms
delay: 5ms
labels:
label: 1
category: socbed-example
B-D Scenario Details for Broad Analysis ConCap
Each scenario evaluates a specific network activity or tool using its default configuration unless noted otherwise. Where applicable, we highlight the use of additional command options. All scenario files, including execution and configuration details, are available in our repository for full reproducibility.
Wfuzz This scenario simulates an HTTP fuzzing attack where the attacker attempts to discover hidden web resources by injecting path variations. The ”common” wfuzz wordlist was used as an additional parameter.
Slowloris This activity performs a Slowloris denial-of-service attack, in which the attacker opens many half-completed HTTP connections to exhaust the server’s resources. The attack was configured to run for 300 seconds, beyond the tool’s default behavior.
Patator FTP This scenario emulates an FTP brute-force attack using Patator. A password list with the 200 most used passwords of 2023 danielmiesslerseclists was explicitly supplied, and failure responses matching the message ”530 Login incorrect.” were ignored using Patator’s ignore option.
Ping A standard ICMP echo test is performed with 10 packets sent to the target host. The number of packets was explicitly specified in the command.
Patator SSH This scenario represents a brute-force attack against an SSH server. Both a username and password list were provided, top usernames shortlist and top 200 passwords from 2023 danielmiesslerseclists , and the failed login attempts are filtered based on the message ”Authentication failed.”
Nmap Version Scan In this scan, Nmap was used with additional options for version detection, SYN scan, and no DNS resolution. The scan targeted ports 79 and 80.
Nmap This scenario runs a more basic Nmap scan, using a SYN scan and disabling DNS resolution. It targeted the same ports (79 and 80) but did not include version detection.
MySQL The attacker connects to a MySQL database and issues a query to read entries from a users table. Default MySQL authentication is used (user: root, password: root), and the server was configured to listen on all interfaces.
Curl/FTP This scenario involves an authenticated file download over FTP. The attacker connects with a predefined username and password to retrieve a specific file (bible.txt). No additional options were used beyond basic authentication and output specification.
Dig (DNS) The attacker sends a series of DNS queries for five popular domain names (e.g., google.com, facebook.com) to the target DNS server on a non-standard port (63). The use of a custom port deviates from the default behavior.
Appendix C Datasets (existing and new ones)
We provide details on the various “datasets” considered in our paper. Specifically, we first provide information on the real-world capture (used in §VI-B), then we describe how we preprocessed the benchmark datasets CICIDS17 and CICIDS18 (used in §VI-A and §VI-C), and finally we provide low-level details of our “new” dataset containing labeled NetFlows related to attacks not contained in currently available benchmarks (mentioned in §VI-C).
C-A Real-word Network Capture: Description
For the experiments in §VI-B, we used data captured in a real-world network representing a “smart home”. Here, we provide more details of this network environment, and the captured PCAP trace and corresponding NetFlow data.
Network Overview The network environment encompasses 40–50 physical devices. Such devices entail: smartphones, laptops / desktops, gaming consoles; as well as various IoT devices (smart speakers, lightbulbs) and media appliances (e.g., smart TV). All these devices are connected to a router through a WiFi 5 or 2.4 interface. The router is connected to the internet through a 50Mbps download speed and 5Mbps upload speed. The devices within the network are kept up-to-date with security patches and their owners are security experts, hence it is safe to assume that the traffic is “benign” (and even if some traffic is “malicious”, it is of a different class than patator meaning that our conclusions are not affected by such circumstances). With regards to the devices used to simulate the attack in this network, they were two laptops both running Ubuntu 18.04; the “target” mounts an Intel i7 7700H with 32GB of RAM; the “defender” mounts an Intel N4100 with 8GB of RAM.
Network Traffic (PCAP). Three sets of network captures were performed, two for benign background traffic and one related to the patator ssh-bruteforce attacks. The background traffic was first captured on the 6th of November 2023 and a second time on the 26th of August 2024, the same day the malicious attacks were executed and captured. The total file size for the benign traffic of the August 26th capture is of 6GB for 8M packets, whereas the attack captures only measured 45MB for 241k packets. For the (benign) trace captured on November 2023, the size is of 17GB for 17M packets
NetFlow. NetFlows were extracted from the network captures using CICFlowMeter. The benign capture from November 2023 contained 30,304 unique NetFlows, while the benign capture from August 2024 contained 49,188 unique NetFlows. The malicious capture, on the other hand, included a total of 7,990 unique NetFlows.
C-B Preprocessing of Benchmark datasets
In this appendix, we provide an overview of the data preprocessing steps undertaken to prepare the network traffic data for analysis. This includes the extraction of NetFlow features using the latest version of CICFlowMeter, data cleaning procedures, and a description of the real world network captures used in our experiments.
NetFlow Extraction The fixed versions of the datasets CICIDS17 and CICIDS18 were released in October 2022. Since then, CICFlowMeter, the tool used to extract the NetFlow features from the network traffic traces, has received over 30 new commits fixing, changing, and adding NetFlow features. To benefit from these updates over the last two years, we replicated the work done by Liu et al. liu2022error by extracting the NetFlows using the current version (Aug ’24) of CICFlowMeter and labeling the flows based on their fixed logic. The “attempted” NetFlows are removed from our dataset.
Data Cleaning. The dataset cleaning steps on the NetFlows performed in our experiments follow the best practices described in previous work apruzzese2022sok ; arp2022and ; d2022establishing . First, the meta-data and spurious features are removed: “id”, “Flow ID”, “Src IP”, “Dst IP”, “Timestamp”, “FWD Init Win Bytes”, and “Bwd Init Win Bytes”. Second, the source and destination ports are mapped to their IANA port categories and subsequently encoded as 0, 1, or 2 for respectively well-known, registered, and dynamic. Third, all NetFlows with missing values are removed. Last, all duplicate NetFlows are removed.
Implementation. The features and hyperparameters used to develop our ML models are provided in our repository.
C-C Generation of (new) CVE Data with ConCap
We generated new labeled ”malicious” data using ConCap by executing three recent CVE-based attack scenarios, ensuring none of them overlap with the datasets mentioned in Flood et al. flood2024bad . These CVEs represent diverse vulnerabilities, each exploited to demonstrate the flexibility of ConCap in handling modern attacks.
  • CVE-2024-47177 cve1 : This vulnerability affects OpenPrinting Cups-Browsed versions 2.0.1 and earlier, where improper handling of the FoomaticRIPCommandLine parameter in PostScript Printer Description (PPD) files allows remote code execution. Attackers can exploit this by creating a malicious IPP (Internet Printing Protocol) server that sends crafted printer information to a vulnerable Cups-Browsed instance, enabling arbitrary command execution on the affected system.
  • CVE-2024-36401 cve2 : In GeoServer versions prior to 2.25.1, 2.24.3, and 2.23.5, unauthenticated remote code execution is possible through unsafely evaluated property name expressions. Specially crafted input can exploit these unsafe evaluations in multiple OGC request parameters, allowing attackers to execute arbitrary code on a vulnerable GeoServer installation.
  • CVE-2024-2961 cve3 : This vulnerability in the GNU C Library (versions 2.39 and earlier) affects the iconv() function, which may overflow the output buffer by up to 4 bytes when converting strings to the ISO-2022-CN-EXT character set. Attackers can exploit arbitrary file read vulnerabilities in PHP applications to escalate to remote code execution by leveraging the iconv() issue, potentially crashing the application or executing code.
The target environments were constructed using Vulhub, an open-source collection of pre-built vulnerable docker environments. The attacker environments were built using official Docker images, combined with the necessary tools and exploit code for each CVE. These environments were then defined into a ConCap scenario together with the attack command and assigned a unique label with the corresponding CVE for automatic traffic labeling. The three scenarios, one for each CVE, are then repeated 320 times in different network environments with a unique set of values for delay, loss, corrupt, and duplicate. Each of these scenarios is repeated for five, two, and four host-space perturbations respectively for CVE-2024-47177, CVE-2024-36401, and CVE-2024-2961.
TABLE VI: Testing existing ML-NIDS against unseen CVEs with ConCap. We develop benchmark ML-NIDS using CICIDS17 (baseline tprtpr=0.999 with fprfpr=0.001) and CICIDS18 (baseline tprtpr=0.999 with fprfpr=0.001). Then, we use ConCap to generate and label NetFlows of three CVEs (we report the # Packets and # NetFlows of each capture) involving completely different attacks than those used to “train” the ML-NIDS, and we test them against all of our models (RF, DT, SVM, DNN, HGB). The results show the average tprtpr (and std. dev.) across all models.
    Traffic Statistics     CICIDS17 CICIDS18
Attack     # Packets # NetFlows     tprtpr (std. dev) tprtpr (std. dev)
CVE-2024-47177     44658 4800     0.205 (0.389) 0.550 (0.396)
CVE-2024-36401     10372 640     0.199 (0.399) 0.214 (0.378)
CVE-2024-2961     391350 1280     0.112 (0.224) 0.011 (0.021)
Appendix D Additional Results and Experiments
We presents detailed results for the interested reader.
First, we provide the training time of the models evaluated on the real-world network in Table VII. Finally, Fig. 6 shows an additional 30 NetFlow features compared for the patator brute force attack between traffic generated by ConCap and on a real network as described in §V.
We provide the tables with the standard deviation of our results (useful to carry out statistical tests) in our repository repository .
D-A Runtime and Computing Results
We report in Table VII the training time of the ML models used for the experiments in §VI. These experiments have been carried out on a machine running Windows 11 Pro (Build 26100) with a Intel 12th Gen Core i7-1265U CPU 1.80 GHz (10 cores, 12 threads) and 32GB RAM.
Then, we report in Table VIII the results of the runtime assessment of ConCap (covered in §VII-A). The laptop used in these tests runs MicroK8s v1.32.3 in a virtualized Linux environment on Windows Subsystem for Linux 2 (WSL2). The host system features an Intel 12th Gen Core i7-1265U CPU 1.80 GHz (10 cores, 12 threads), 32GB RAM, and Windows 11 Pro (Build 26100), WSL2 has access to 12 logical CPUs and 23 GiB of RAM. During our tests, 12 variations of nmap scenarios are executed in a single ConCap run configured to run three scenarios concurrently. ConCap shows minimal differences in initialization time and resource usage across the two setups. Interestingly, initialization time is slightly lower on the laptop despite its more limited resources. We attribute this to the elimination of network communication delays between the machine running ConCap and the cluster executing the scenarios, as the microK8s setup is entirely local. Initialization time includes spawning the attacker and target containers, configuring the scenario network environment, and activating packet capture on the targets, as well as associated network delays. We also measured the resource consumption of the processing pods: both argus and cicflowmeter. While argus remains lightweight, cicflowmeter consistently exhibits an order of magnitude higher CPU and memory usage compared to the scanning scenarios themselves—highlighting the relative cost of traffic processing versus generation. Although both environments perform similarly, the Kubernetes cluster provides horizontal scalability, enabling the execution of an increased number of scenarios in parallel when needed.
TABLE VII: Training times for the models for the assessment of the real-world network. Avg time (sec) and std. dev. over 5 training folds.
Train Set Benign + P=1 Benign + P=0
DT $0.3\pm 0.0$.0) $0.2\pm 0.0$.0)
RF $3.1\pm 0.0$.1) $4.1\pm 0.0$.2)
XGB $3.0\pm 0.0$.4) $3.9\pm 0.0$.6)
SVM $0.5\pm 0.0$.0) $0.9\pm 0.0$.1)
DNN $9.2\pm 0.0$.5) $12.3\pm 0.0$.5)
(a) Real-world
Train time
DT $6.7\pm 0.1$.3)
RF $40.8\pm 0.4$.2)
XGB $15.1\pm 0.3$.8)
SVM $4.1\pm 0.0$.9)
DNN $17.2\pm 0.2$.0)
(b) CICIDS17
Train time
DT $40.4\pm 0.1$.8)
RF $547.8\pm 0.6$.3)
XGB $118.8\pm 0.7$.4)
SVM $101.3\pm 1.0$.9)
DNN $574.5\pm 4.5$.8)
(c) CICIDS18
TABLE VIII: Performance evaluation of ConCap running various port scan scenarios on a multi-node K8s cluster versus microK8s on a WSL-enabled laptop. The results show minimal differences in initialization time and resource usage between the two environments. However, the K8s cluster offers horizontal scalability, allowing for greater resource availability and enabling more parallel scenario executions.
Cluster Laptop
Scenario Init Time (s) CPU (mcpu) Mem (MiB) Init Time (s) CPU (mcpu) Mem (MiB)
nmap-pn-$s$ 2.202 4.10 1.67 2.300 2.77 0.68
nmap-pn-$s$-sv 2.696 20.94 1.67 2.155 2.67 2.00
nmap-pn-st 2.622 7.89 1.67 1.866 4.37 1.63
nmap-pn-st-sv 2.737 3.33 1.67 2.189 2.53 2.00
nmap-pn-su 2.747 3.33 1.67 2.095 4.52 1.67
nmap-pn-su-sv 2.724 19.76 1.67 2.869 1.00 1.67
nmap-$s$ 2.679 3.00 1.67 2.300 1.00 1.67
nmap-$s$-sv 4.682 10.88 1.67 2.155 11.03 2.00
nmap-st 2.286 4.67 1.67 1.866 4.40 1.67
nmap-st-sv 3.996 15.17 1.67 2.189 8.71 2.00
nmap-su 3.767 8.00 1.67 2.095 3.00 1.67
nmap-su-sv 2.536 1.00 1.67 2.869 8.00 1.67
argus 18.19 0.00 14.88 0.00
cicflowmeter 119.85 2.98 109.10 10.95
ConCap 33.18 35.29 29.95 33.76
Avg Init Time 2.97 s 2.25 s
Figure 5: Using ConCap to reproduce complex attack chains envisioned in MITRE ATT&CK. The attacker first compromises an exposed target before performing lateral movement to two internal hosts.
D-B Reproducing Multi-step and Multi-host attacks
We conclude our practical demonstrations by showing how to use ConCap to carry out experiments entailing sophisticated attacks. Specifically, we demonstrate that ConCap enables automatic reproduction of multi-host and multi-step attacks.
Threat Model. We assume a network consisting of four or more hosts, each having different privileges and containing different information. An attacker has obtained control of one host, and wants to expand their control and steal sensitive information. However, the attacker does not know the network topology, and must hence carry out reconnaissance activities to identify vulnerabilities that can be exploited to reach the intended goal. In other words, we envision an attack chain spanning multiple stages and hosts, which can be mapped to a range MITRE ATT&CK tactics mitreattack . Altogether, the attack encompasses the following stages (expressed via the well-known MITRE ATT&CK terminology): Gather Victim Network Information, Network Service Discovery, Brute Force, Internal Proxy, SSH Lateral Movement, Data Collection from Local System, Exfiltration Over C2, and Exploitation for Privilege Escalation. Our intention is showing how to use ConCap to generate (and automatically label) NetFlows of the entire attack chain—in an end-to-end fashion.
Implementation. To replicate the setting envisioned in our threat model, we defined the scenario in Listing 5 (in the Appendix). Such a scenario embeds the use case, schematically depicted in Fig. 5, in which an attacker first performs a port scan against an SSH server using nmap nmap , then launches a brute-force SSH attack using Medusa medusa . Upon discovering valid credentials, the attacker uses SSH to set up a SOCKS proxy through the compromised host and scans additional internal hosts (a MySQL database and a vulnerable libssh server). Then, the attacker forwards internal ports via SSH tunneling to the local machine, allowing authenticated access and data exfiltration against the MySQL target and enabling the attacker to exploit a known vulnerability against libssh.
Considerations. In practice, when ConCap processes such a scenario, all of the aforementioned operations are automatically carried out. We also note that each target in this scenario is separately labeled with its function (e.g., “brute-force-ssh”, “exfiltration”, or “cve”) and configured with specific network constraints (e.g., bandwidth and delay). At the end of the entire experiment, the OpenSSH server received 3635 network packets corresponding to 184 CICFlowMeter and 185 argus NetFlows. The MySQL server received 280 packets corresponding to 109 CICFlowMeter and 110 argus NetFlows. The libSSH server received 246 packets corresponding to 106 CICFlowMeter and 106 argus NetFlows. All NetFlows (which we provide in our repository repository ) are labeled according to the specifications of the scenario. Notably, by sharing this scenario (i.e., Listing 5), future researchers can inspect the entire traffic-generation process: this is important for open science and to address the skepticism around NIDS research flood2024bad .
D-C Testing additional ML-based NIDS with ConCap
In our paper (§VI), we showed that ConCap can be used to test “simple” ML-based (e.g., RF or SVM) NIDS. Here, we expand our assessment by showing that ConCap’s generated data can be used also by other families of ML-driven NIDS.
Please note that our goal is merely to show that ConCap’s data can be used to train and test such additional methods: we do not seek to (a) propose new methods, (b) outperform/tweak existing ones, or (c) benchmark prior work.
Considered models. There are hundreds of papers proposing, or evaluating, methods to detect network intrusions via ML apruzzese2023sok . For the sake of this additional demonstration, we consider the methods considered in the recent USENIX’23 paper xNIDS wei2023xnids . Specifically, this work considers two families of deep-learning–based methods for intrusion detection: one reliant on AutoEncoders (drawn from mirsky2018kitsune ; shone2018deep ), which we denote as AE-IDS; and another one reliant on Recurrent Neural Networks (drawn from yin2017deep ; jan2020throwing ), which we denote as RNN-IDS. Note that, altogether, the papers proposing these methods have thousands of citations on Google Scholar (as of December 2025) or are published in top-tier venues (e.g., S&P’22, NDSS’18). Then, to provide yet-another perspective, we also consider transformer-based approaches which empower multimodal large-language models (LLMs), inspired by the recent luay2025multimodal .
Implementation and Datasets. We take the code from wei2023xnids (which is publicly available) as a blueprint for our AE-IDS and RNN-IDS models; whereas, for LLMs, we consider: ChatGPT 5.1, and Gemini Flash 2.5 (both of which being the best freely-available commercial models as of December 2025). For both of these LLMs, inspired by luay2025multimodal , we will assess them with a zero-shot prompting strategy (with the prompt in Listing 7), as well as in an “zero-shot-with-augmentation” fashion (with the prompt in Listing 8). [^6] Note that the approach in luay2025multimodal uses two-dimensional plots as input to the LLM, whereas we use ConCap’s raw data. This is because we want to show that ConCap can be used off-the-shelf to test LLMs. However, there is nothing preventing one from generating the same two-dimensional plots of luay2025multimodal (which simply show the source and destination IP for the x- and y-axes, and points indicate the amount of exchanged bytes—all of which being metrics derivable from ConCap’s data. For the evaluation dataset, we consider the same used in §VI-A, i.e., a mix of CICIDS17 and the data generated by ConCap referring to the ssh-patator attack. Such a setup is valid for the sake of our demonstration, given that the approaches we considered have been evaluated in completely different setups (e.g., the “outdated” NSL-KDD or the CIC-DoS2017 jazi2017detecting ).
Evaluation and Results. We replicate the experiments in §VI-A. Specifically, we train each model (AE-IDS and RNN-IDS) on 80% of the benign data in CICIDS17, as well as on 80% of the data generated by ConCap for ssh-patator (launched with default options); and then we test it on the remaining 20% (benign and malicious). We also consider doing a mixed experiment by testing each model on 20% of the ssh-patator included in CICIDS17; as well as by creating another variant trained on 80% of the ssh-patator included in CICIDS17 and testing it on 20% of the ssh-patator generated by ConCap. For the LLMs, the zero-shot prompt does not have any training (barring that to develop the LLM itself), whereas the augmented-zero-shot prompt embeds a fine-tuning step done on the same data used to train the AE-IDS and RNN-IDS. We repeat our experiments five times to ensure consistency. Let us discuss the results of our experiments:
  • AE-IDS: This experiment confirms our conclusions drawn from §VI-A. Specifically, when trained [^7] Training the AE-IDS required $\approx$10 minutes on our system. on ConCap’s generated data (which is always malicious), the AE-IDS achieves tprtpr=0.999 on the testing partition of ConCap’s generated data, and tprtpr=0.987 on the testing partition of malicious data of the same attack included in CICIDS17; conversely, when trained on the malicious data of CICIDS17, the AE-IDS achieves tprtpr=0.955 on the testing partition of CICIDS17, and tprtpr=0.952 on the testing partition of ConCap’s generated data. The fprfpr is 0.478 on the former case, and 0.209 in the latter case: such an underwhelming result is because our AE-IDS uses the same thresholding mechanism used in xNIDS wei2023xnids (note that AutoEncoders are not classifiers: to use them in a classification task, one must specify a threshold on the reconstruction error. We used the one of xNIDS, which was derived on a different dataset. To improve the fprfpr, one can simply change the threshold). The different fprfpr is because the number of datapoints in the training set is different across the two experiments, but what is crucial is that, for this detector, the data generated by ConCap is essentially equivalent to that of CICIDS17, since the classification results on these two groups are statistically the same (verified with a t-test: $p<.05$).
  • RNN-IDS: For this experiment, we report that, after 2 hours of training, the underlying RNN did not reach convergence. Therefore, given our previous experiment, as well as those in our main paper (§VI-A) showing that simple models (such as RF or SVM) can achieve near-perfect detection accuracy while requiring seconds to train (see Table VII(b)), we conclude that using RNN-IDS on our dataset is not a wise choice and stopped this experiment. However, what matters is that the RNN-IDS was being trained, meaning that ConCap’s data was “compatible” with that expected in RNN-IDS, thereby validating the goal of these experiments (note that SaTML’26 call for papers explicitly states to be mindful of the computing resources spent in training ML models, and we believe that additional training would not have changed our conclusions).
  • ChatGPT: When assessed in the “zero-shot” prompt, the model always classified each datapoint as benign. This is expected: the malicious datapoints are just repeated SSH attempts, which not necessarily imply malicious intent. However, by using the “augmented-zero-shot” prompt (which uses ConCap’s generated labelled data to “teach” the LLM that quick bursts of SSH connections are malicious), the model achieves a perfect classification accuracy (i.e., tprtpr=1.000 and fprfpr=0.000). We inspected the output provided by ChatGPT, and found that, to derive such a result, the model learned heuristics that enabled a clear separation of the benign from malicious datapoints. (note that these results align with those in luay2025multimodal ).
  • Gemini: The results for Gemini 2.5 match those of ChatGPT.
The code of these experiments is in our repository repository .
Takeaway. Our experiments show that ConCap can be used to develop and assess multiple families of ML-based NIDS. Moreover, our results further confirm that ConCap’s generated data can be functionally equivalent to that generated in other environments (e.g., that of CICIDS17).
Listing 7: Zero-shot prompt used to evaluate LLMs with ConCap’s data.
You are a cybersecurity analyst working in a Security Operations Center (SOC). You are analyzing a CSV file that contains NetFlow records.
Data Description:
Each row in the CSV represents a network flow with features such as: Src Port, Dst Port, Protocol, Flow Duration, Number of packets Your task is to classify each row as either: 0 = Benign 1 = Malicious
Your Tasks:
Determine whether the NetFlow record shows suspicious behavior. Output only 0 or 1 for each row (0 = benign, 1 = malicious) on a separate row and write to csv file.
Listing 8: Augmented zero-shot prompt used to evaluate LLMs with ConCap’s data. We augment the prompt in Listing 7 by instructing the LLM to use labeled data (in the X_train.csv file) as a guide for its inference.
You are a cybersecurity analyst working in a Security Operations Center (SOC). You are analyzing a CSV file that contains NetFlow records (X_test_shuffled.csv).
Data Description:
Each row in the CSV represents a network flow with features such as: Src Port, Dst Port, Protocol, Flow Duration, Number of packets Your task is to classify each row as either: 0 = Benign 1 = Malicious
Your Tasks:
Determine whether the NetFlow record shows suspicious behavior. Base your predictions on the provided labeled NetFlows (X_train.csv and y_train.csv). Output only 0 or 1 for each row (0 = benign, 1 = malicious) on a separate row and write to csv file.
Ack Flag Count
Average Packet Size
Bwd Act Data Packets
Bwd Header Length
Bwd IAT Mean
Bwd IAT Std
Bwd IAT Total
Bwd PSH Flags
Bwd Packets/s
Down/Up Ratio
Flow Bytes/s
Flow Duration
Flow IAT Mean
Flow IAT Std
Flow Packets/s
Fwd Act Data Packets
Fwd Header Length
Fwd IAT Mean
Fwd IAT Std
Fwd IAT Total
Fwd Packets/s
Idle Mean
PSH Flag Count
Packet Length Mean
Packet Length Std
Total Bwd Packets
Total Connection Flow Time
Total Fwd Packet
Total Length of Bwd Packet
Total Length of Fwd Packet
Figure 6: Comparison of NetFlow feature distributions of a Patator brute-force SSH attack between ConCap and a real network.